### v3

# On the jump of the cover time in random geometric graphs

Abstract.
In this paper we study the cover time of the simple random walk on the giant component of supercritical $d$-dimensional random geometric graphs on $\mathrm{Poi}(n)$ vertices. We show that the cover time undergoes a jump at the connectivity threshold radius $r_{c}$: with $r_{g}$ denoting the threshold for having a giant component, we show that if the radius $r$ satisfies $(1+\varepsilon)r_{g}\leq r\leq(1-\varepsilon)r_{c}$ for $\varepsilon>0$ arbitrarily small, the cover time of the giant component is asymptotically almost surely $\Theta(n\log^{2}n$). On the other hand, we show that for $r\geq(1+\varepsilon)r_{c}$, the cover time of the graph is asymptotically almost surely $\Theta(n\log n)$ (which was known for $d=2$ only for a radius larger by a constant factor). Our proofs also shed some light onto the behavior around $r_{c}$.
Dieter Mitsche has been partially supported by grant GrHyDy ANR-20-CE40-0002 and by Fondecyt grant 1220174.
Keywords: cover time, random walk, random geometric graphs
MSC Class: 05C80, 60D05, 05C81
1. Introduction
We denote by $\mathcal{G}(n,r,d)$ the by now classic model of random geometric graphs defined as follows: let $d$ and $n$ be positive integers, and let $r$ be a positive real number, possibly depending on $n$. The vertices of the graph are the points of a Poisson Point Process (PPP) of intensity $1$ in $\Lambda_{n}=\left[-\frac{n^{1/d}}{2},\frac{n^{1/d}}{2}\right]^{d}$, that is, the $d$-dimensional axis-parallel cube of volume $n$ centered at the origin, and two vertices are connected by an edge if their Euclidean distance is at most $r$. Random geometric graphs were originally introduced by Gilbert [24] as a model for telecommunication networks, and since then they received a lot of attention both from a theoretical point of view (see for example the monograph of Penrose [35], or also [36]) as well as from an applied point of view (see for example [3, 33, 23]).
The simple random walk on the vertices of a (throughout the paper finite and connected) graph $G=(V,E)$ is a discrete-time stochastic process with random variables $X_{0},X_{1},\ldots$ such that $X_{0}=v$ for some $v\in V$, and for $i\geq 0$, $X_{i+1}$ is a vertex chosen uniformly at random from the neighbors of $X_{i}$. For a graph $G(V,E)$, the cover time of the simple random walk is then defined as
$\tau_{cov}(G)=\max_{v\in V,X_{0}=v}\mathbb{E}[\min\{t\geq 0:\bigcup_{s=0}^{t}\{X_{s}\}=V\}].$
In words, the cover time is the expected time it takes to visit all vertices, when starting from a worst-case vertex.
1.1. Main results
In order to state our results, we need a bit of additional notation. For a sequence $\{\mathcal{A}_{n}\}_{n}$ of events, we say that $\mathcal{A}_{n}$ holds asymptotically almost surely (a.a.s.), if $\Pr[\mathcal{A}_{n}]\to 1$, as $n\to\infty$. For $G\in\mathcal{G}(n,r,d)$ (identifying $G$ with its vertex set), define $r_{c}=r_{c}(n,d)$ as the smallest radius such that a.a.s. a random geometric graph with such a radius is connected. It is well known (see [25]) that $r_{c}=(\log n/V_{d})^{1/d}$, where $V_{d}$ is the volume of the unit ball in $d$ dimensions. Define also $r_{g}=r_{g}(n,d)$ as the smallest radius such that a.a.s. a random geometric graph with such a radius has a giant component, that is, a component of size linear in $n$. It is well known that $r_{g}=\Theta(1)$ (the existence of such a threshold is well known, the exact value is only conjectured, even in $2$ dimensions). For rigorous lower and upper bounds on $r_{g}$ see [35].
The cover time of $\mathcal{G}(n,r,d)$ was, as far as we know, only studied in two papers: first, Avin and Ercal (see [5]) showed that for $G\in\mathcal{G}(n,r,2)$ with $r^{2}>8\log n$, that is, a constant factor above the connectivity threshold, $\tau_{cov}(G)=\Theta(n\log n)$ a.a.s. Second, Cooper and Frieze [12] showed that for $d\geq 3$ and $G\in\mathcal{G}(n,r,d)$ with $r=(c\log n/V_{d})^{1/d}$, with $c>1$ being arbitrary and $V_{d}$ as before being the volume of the $d$-dimensional unit ball, a.a.s. $\tau_{cov}=c\log\left(\frac{c}{c-1}\right)n\log n$ (their result does not apply for $d=2$). In this paper, we look at sparser graphs not dealt with by these papers: for $G\in\mathcal{G}(n,r,d)$ and $r\geq r_{g}$, we denote by $L_{1}(G)$ its largest component (that is, the largest connected graph induced by the vertex set of this component). Our first main result is the following:
Theorem 1.1.
Let $\varepsilon>0$ be arbitrarily small. Let $d\geq 2$. Let $G\in\mathcal{G}(n,r,d)$ be a random geometric graph with $(1+\varepsilon)r_{g}\leq r\leq(1-\varepsilon)r_{c}$. Then, a.a.s., $\tau_{cov}(L_{1}(G))=\Theta(n\log^{2}n)$.
Our second main result complements the previous result, by considering a random geometric graph right above the connectivity threshold (which was left open in the case $d=2$):
Theorem 1.2.
Let $\varepsilon>0$ be an arbitrarily small constant. Let $d=2$. Let $G\in\mathcal{G}(n,r,d)$ be a random geometric graph with $r\geq(1+\varepsilon)r_{c}$. Then, a.a.s.  $\tau_{cov}(G)=\Theta(n\log n)$.
Theorem 1.1 and 1.2, together with known results for $d\geq 3$, show that for $G\in\mathcal{G}(n,r,d)$, $\tau_{cov}$ undergoes a jump at $r_{c}$, decreasing from $\Theta(n\log^{2}n)$ to $\Theta(n\log n)$. Our proofs shed some light onto the behavior of $\tau_{cov}$ right at $r_{c}$, and we comment on this in the concluding remarks (we do not know the exact behavior, though).
1.2. Further related work
Apart from the aforementioned papers, the cover time is a classic parameter in the study of random walks that has been analyzed in many contexts and on different graphs: on the one hand, for general graphs, Feige showed in [19, 20] that for any connected graph on $n$ vertices, the cover time is a.a.s. at least $(1-o(1))n\log n$ and at most $(4/27)n^{3}+o(n^{3})$, and he also gave graphs attaining the asymptotic bounds.
On the one hand, from an algorithmic point of view, for general graphs, Kahn, Kim, Lóvasz and Vu in [26] first established a $(\log\log n)^{2}$-polynomial time approximation algorithm for the cover time. Later this was improved by Ding, Lee and Peres, who in a breakthrough result in [18] found a polynomial-time approximation algorithm approximating the cover time up to a constant factor (later this result was refined by Ding [17] and Zhai [38]).
Again in another context, on the random graph $\mathcal{G}(n,p)$, a series of papers of Cooper and Frieze analyzed the cover time of the largest component of $\mathcal{G}(n,p)$ for different values of $p$ (summarized in Theorem 1 of [22]). Moreover, in the aforementioned paper [22], Frieze, Pegden and Tkocz obtained more precise estimates for the cover time of the emerging giant component, improving on previous results of Barlow, Ding, Nachmias and Peres [6] and Cooper, Frieze, Lubetzky [13].
Finally, in a geometric setup that is the most similar to the current work, in a series of papers, the cover time of the $d$-dimensional toroidal grid of side length $n$ (and thus with $n^{d}$ vertices) has been studied, and remarkably precise results are known by now. For $d=2$, an upper bound of the cover time of the correct order was first found by Aldous [4], and a non-matching first order lower bound by Lawler [29]. Finally, the correct first order was obtained in the landmark paper of Dembo, Peres, Rosen and Zeitouni [15]: they showed that the cover time of the $2$-dimensional torus satisfies
$\tau_{cov}/(n^{2}\log^{2}n)\to 4/\pi,$
as $n\to\infty$, where the convergence is in probability. Ding [16] improved the result by showing that $\sqrt{\tau_{cov}/n^{2}}-\frac{2}{\sqrt{\pi}}\log n$ is of order $\log\log n$ with probability tending to $1$ as $n\to\infty$. Abe [2] then found the second order term: he showed that there exists $c\in(0,1)$, such that
$2\log n-\log\log n-(\log\log n)^{c}\leq\frac{\tau_{cov}}{\frac{2}{\pi}n^{2}\log n}\leq 2\log n-\log\log n+(\log\log n)^{c}$
with probability tending to $1$ as $n\to\infty$, similar to a previous bound in the continuous setup of Brownian motion by Belius and Kistler [8]. Recently, using a comparison with the extremal landscape of the discrete Gaussian free field, Louidor and Saglietti [32] went even further: they showed that the square root of the cover time normalized by the size of the subset is tight around $\frac{1}{\sqrt{\pi}}\log n-\frac{1}{4\sqrt{\pi}}\log\log n$. The latter approach was inspired by an analysis of the cover time on the binary tree of depth $n$: in this case, Cortines, Louidor and Saglietti showed in [14] that the cover time, normalized by $2^{n+1}n$, and then centered by $(\log 2)n-\log n$, admits a weak limit as $n\to\infty$. The limiting distribution in this case obtained was a Gumbel random variable with rate one, shifted randomly by the logarithm of the sum of the limits of the derivative martingales associated with two negatively correlated discrete Gaussian free fields on the infinite version of the tree. For $d\geq 3$, Belius in [7] showed that, normalized by the volume of the torus and then centered by a constant times $\log n^{d}$, the cover time converges in law to a Gumbel distribution.
In the context of a percolated grid (of side length $n$), some results are known as well: whereas the cover time of (unpercolated) $\mathbb{Z}^{2}$ is a.a.s. of the order $n^{2}\log^{2}n$ and of (unpercolated) $\mathbb{Z}^{d}$ for $d\geq 3$ of the order $n^{d}\log n$, Abe in [1] showed that for any $d\geq 2$, the cover time of the giant component of supercritical Bernoulli percolation in $\mathbb{Z}^{d}$ is a.a.s. of the order $n^{2}\log^{2}n$, thereby showing a quantitative difference of unpercolated and percolated $\mathbb{Z}^{d}$ for $d\geq 3$.
In the context of random hyperbolic graphs, recently, Kiwi, Schepers and Sylvester studied the cover time of the giant component of random hyperbolic graphs on $n$ vertices. They showed that the cover time of the giant component of these graphs is $\Theta(n\log^{2}n)$ (see [28]).
1.3. Outline of proofs
The approach we follow is based on the idea of seeing the graph as an electric network. The main focus of the paper is therefore to obtain good enough bounds on the effective resistance.
The lower bound of the proof of Theorem 1.1 is comparably simpler: depending on the regime of the radius, either we find pending paths of order $\log n/r^{d}$ (as in the paper of Abe [1]), or we will find many vertices of degree $1$ inside the giant component. The upper bound is also inspired by Abe’s result in [1] on $\tau_{cov}$ of a supercritical percolation cluster on $\mathbb{Z}^{d}$: we first mimic Abe’s backbone construction from [1] over Bernoulli percolation now using a suitable renormalization scheme on the random geometric graph for fixed radius $r$. We then use this structure to build a flow that dissipates energy in a convenient way: we have to be careful with the construction as there are additional obstacles that do not appear in Bernoulli percolation, such as the fact that crossings could be arbitrarily close to one another or that we may have to extract induced paths from paths in densely populated areas. When the typical degree is tending to infinity, one extra splitting of the flow is introduced in order to bound the effective resistance between any two vertices of a certain backbone (which is large and well spread out) by $O(\log n/r^{2d})$.
The proof of Theorem 1.2 elaborates on the idea of the upper bound of Theorem 1.1 by an additional ingredient: on the one hand, since the typical degree is high, we need again the extra splitting from the case when the radius tends to infinity. On the other hand, we then show that other vertices outside this backbone cannot be separated from the backbone by taking out too few vertices: we will find many short internally vertex-disjoint paths between any vertex outside and some vertex inside the backbone, on which we can split the flow again evenly.
Further notation. For $\mathrm{j}\in\mathbb{Z}^{d}$, we denote by $\mathrm{j}_{i}$ the $i$-th coordinate. We denote by $e_{i}$ the $i$-th standard basis vector in $d$ dimensions. For a set $A$, we denote its cardinality by $|A|$. For $G=(V,E)$ and $x,y\in V$, we write $x\sim y$ for $\{x,y\}\in E$. We denote by $d_{G}(x,y)$ the graph distance between $x$ and $y$, that is, the number of edges on a shortest path between $x$ and $y$. If $x$ and $y$ are identified with points, we also denote by $d_{E}(x,y)$ the Euclidean distance between $x$ and $y$. We say that two paths of vertices $u_{0}:=x,u_{1},\ldots,u_{m},u_{m+1}:=y$ and $v_{0}:=x,v_{1},\ldots,v_{n},v_{n+1}:=y$ between $x$ and $y$ are internally vertex-disjoint, if $\{u_{0},\ldots,u_{m+1}\}\cap\{v_{0},\ldots,v_{n+1}\}=\{x,y\}$. A path of vertices $v_{1},\ldots,v_{k}$ in $G(V,E)$ is induced, if $\{v_{i},v_{j}\}\in E$ implies $j=i+1$ for some $i\in\{1,\ldots,k-1\}$. Note that in an induced path of $G\in\mathcal{G}(n,r,d)$ for every vertex of the path except for the first and last there are exactly two other vertices in the path at Euclidean distance at most $r$, and for the first and last vertex there is only one other vertex in the path at Euclidean distance at most $r$. Observe also, by taking out vertices if needed, one may obtain an induced path from a path that is not, and that has the same endpoints as the original path.
Plan of paper. In Section 2 we recall basic facts on random walks, in particular on the effective resistance, as well as facts on random geometric graphs and Bernoulli percolation. In Section 3 we then introduce a renormalization scheme that will turn out to be very useful in our proofs. Section 4 is then dedicated to the proof of the lower bound of Theorem 1.1. In Section 5 we then show the corresponding upper bound of Theorem 1.1 for $r=\Theta(1)$. In Section 6 we extend this to the case $r=\omega(1)$. Finally, in Section 7 we prove Theorem 1.2.
2. Preliminaries
In this section we recall basics on different topics, and we also prove some easy preliminary results.
2.1. Markov chains and random walks
A network is a connected graph $G=(V,E)$ together with a set of positive real values $(\mathfrak{r}(e))_{e\in E}$, the so called resistances of the edges. We denote the network simply as $G$ and make no distinction with the underlying graph. We define the conductance $\mathfrak{c}$ of the edge $e$ as $\mathfrak{c}(e)=\frac{1}{\mathfrak{r}(e)}$. Given a network, we define the associated transition probabilities by
$P(x,y)=\frac{\mathfrak{c}({x,y})}{\mathfrak{c}(x)}$ (2.1)
where $\mathfrak{c}(x)=\sum_{y\sim x}\mathfrak{c}(x,y)$. For $\{x,y\}\in E$, denote by $\overrightarrow{e}=\overrightarrow{xy}$ an oriented edge, oriented from $x$ to $y$, and by $-\overrightarrow{e}=\overrightarrow{yx}$ the oriented edge, oriented from $y$ to $x$. For $x,y\in V$, a unit flow from $x$ to $y$ is a function $\theta$ on oriented edges such that $\theta(\overrightarrow{xy})=-\theta(\overrightarrow{yx})$, $\sum_{x\sim a}\theta(\overrightarrow{xa})=1$, $\sum_{a\sim y}\theta(\overrightarrow{ay})=1$ and
$\sum_{z\sim a}\theta(\overrightarrow{za})=0\text{ for all $z\in V\setminus\{x,y\}$.}$ (2.2)
Next, we have a central definition:
Definition 2.1.
The energy of a unit flow $\theta$ is given by
$E(\theta)=\sum_{e\in E}\theta^{2}(e)\mathfrak{r}(e).$ (2.3)
The effective resistance between $x,y\in V$ is then defined as
$\mathcal{R}(x\longleftrightarrow y,G)=\inf\{E(\theta):\text{ $\theta$ is a unit flow from $x$ to $y$}\}$ (2.4)
Remark 2.2.
By definition, we may get upper bounds on the effective resistance by specifying a unit flow between $x$ and $y$. A trivial upper bound for $\mathcal{R}(x\longleftrightarrow y,G)$ is therefore given by the graph distance between the vertices (define the unit flow over the shortest path with all edges having flow equal to $1$).
The effective resistance can be interpreted as the value of the resistance that appears when replacing the entire network with only one edge between $x$ and $y$ that mimics the resistance between them in the original network. A sub-network $G^{\prime}$ of $G$ is a connected sub-graph of $G$ with the same values of resistances on the remaining edges. We now collect a few results that will turn very useful for us. First, effective resistance can not decrease when we delete edges:
Theorem 2.3 (Rayleigh, Theorem 9.12 of [30]).
Let $G^{\prime}$ be a sub-network of $G$. Then, for $x,y\in G^{\prime}$ we have
$\mathcal{R}(x\longleftrightarrow y,G)\leq\mathcal{R}(x\longleftrightarrow y,G^{\prime}).$ (2.5)
When there is no room for confusion on the underlying network $G$ we simply denote the effective resistance by $\mathcal{R}(x\longleftrightarrow y)$. Next, we state that the triangle inequality holds for the effective resistance:
Lemma 2.4 (Corollary 10.8 of [30]).
Let $G=(V,E)$ and let $x,y,z\in V$. Then
$\mathcal{R}(x\longleftrightarrow z)\leq\mathcal{R}(x\longleftrightarrow y)+\mathcal{R}(y\longleftrightarrow z).$ (2.6)
For $G=(V,E)$ and $x,y\in V$, we say that $\Pi\subseteq E$ is an edge-cutset separating $x$ from $y$ if every path from $x$ to $y$ contains at least one edge in $\Pi$. In other words, if we remove $\Pi$ from $E$ then $x$ is isolated from $y$. We will use the following Nash-Williams inequality to obtain lower bounds for effective resistances.
Lemma 2.5 (Proposition 9.15 of [30]).
Let $\{\Pi_{k}\}_{k}$ be a pairwise edge-disjoint family of edge-cutsets that separate $x$ from $y$. Then
$\mathcal{R}(x\longleftrightarrow y)\geq\sum_{k}\left(\sum_{e\in\Pi_{k}}\mathfrak{c}(e)\right)^{-1}.$ (2.7)
Note that for $G=(V,E)$, the simple random walk admits a natural network representation when the set of resistances is given by $\mathfrak{r}(e)=1$ for all $e\in E$. We next relate $\tau_{cov}$ to the effective resistance:
Theorem 2.6 (Theorem 2.4 of [10]).
Let $G=(V,E)$ be a connected graph. Then there exists a universal constant $\gamma_{1}>0$ such that
$\tau_{cov}\leq\gamma_{1}|E|\log(|V|)\left(\max_{x,y\in V}\mathcal{R}(x\longleftrightarrow y)\right).$ (2.8)
Theorem 2.7 (Theorem 1.3 of [26]).
Let $G=(V,E)$ be a connected graph. Given $V^{\prime}\subseteq V$, there exists a universal constant $\gamma_{2}$ such that
$\tau_{cov}\geq\gamma_{2}|E|\log(|V^{\prime}|)\left(\min_{\begin{subarray}{c}x\neq y\\ x,y\in V^{\prime}\end{subarray}}\mathcal{R}(x\longleftrightarrow y)\right).$ (2.9)
2.2. Random geometric graphs
We next state and prove a few useful facts about random geometric graphs. We start with a large deviation bound for the number of edges (the result is much more general, we only state a simplified form that is sufficient to deduce the corollary needed for our purposes): We recall that a sequence of random variables $X_{n}$ satisfies an upper tail large deviation principle with speed $s(n)$ and rate function $I(x)$, if for any closed set $F\subset[0,\infty)$,
$\limsup_{n\to\infty}\frac{1}{s(n)}\log\mathbb{P}\left(\frac{X_{n}-\mathbb{E}[X_{n}]}{\mathbb{E}[X_{n}]}\in F\right)\leq-\inf_{x\in F}I(x),$
and for any open set $G\subset[0,\infty)$,
$\liminf_{n\to\infty}\frac{1}{s(n)}\log\mathbb{P}\left(\frac{X_{n}-\mathbb{E}[X_{n}]}{\mathbb{E}[X_{n}]}\in G\right)\geq-\inf_{x\in G}I(x).$
Theorem 2.8 (Theorem 2.2 of [11]).
Denote by $V_{d}$ the volume of the $d$-dimensional unit ball. Let $G=(V,E)\in\mathcal{G}(n,r,d)$ with $r=\Theta(1)$, so that $\mu_{n}:=\mathbb{E}[|E|]=V_{d}nr^{d}/2$. Let $I(x):=\left(\frac{1}{2}\right)\sqrt{2x}$. Then $|E|$ satisfies an upper tail large deviation principle with speed $\sqrt{\mu_{n}}\log n$ and rate function $I(x)$.
We obtain the following corollary:
Corollary 2.9.
Let $d\geq 2$, and let $G\in\mathcal{G}(n,r,d)$ with $r=\Theta(1)$. A.a.s., every axis-parallelepiped with integer coordinates (inside $\Lambda_{n}$) of side lengths $Mr\log n$ for some constant $M>0$ in 2 dimensions and $Mr$ in the remaining $d-2$ dimensions contains $O(\log^{2}n)$ vertices and has $O(\log^{2}n)$ edges with at least one endpoint in it.
Proof.
First note that by standard Chernoff bounds for Poisson point processes, together with a union bound over all $O(n)$ parallelepipeds, each such parallelepiped contains a.a.s. $O(\log^{2}n)$ vertices, showing the first part of the corollary. Condition on this below. Fix then any parallelepiped $P$ and denote by $X_{v}$ the degree of vertex $v$. Observe that the expected degree of each vertex $v$ then satisfies $\mathbb{E}[X_{v}]=\Theta(1)$, and the total expected degree of vertices $X=\sum_{v\in P}X_{v}$ in $P$ satisfies $\mathbb{E}[X]=O(\log^{2}n)$. Thus, the same asymptotic bound holds for the number of edges $|E|$ with at least one endpoint in $P$. By Theorem 2.8, for any constant $c>0$, $\mathbb{P}((|E|-\mathbb{E}[|E|])\geq c\mathbb{E}|E|)\leq e^{-I(c)\sqrt{\mu_{n}}\log n}$. For every $M$, since $I(c)$ is increasing in $c$, we may thus choose $c=c(M)$ large enough, so that $I(c)\sqrt{\mu_{n}}\geq 1.1$. Thus, with probability at least $1-n^{-1.1}$, inside $P$ the number of edges is $O(\log^{2}n)$, and by a union bound over all $O(n)$ parallelepipeds, the corollary follows. ∎
The following standard result will be used several times in the sequel. Since we were unable to find a reference, we give a proof for completeness.
Lemma 2.10.
Let $d\geq 2$, let $G\in\mathcal{G}(n,r,d)$, and let $r\geq(1+\varepsilon)r_{g}$ for arbitrarily small $\varepsilon>0$. A.a.s., the number of edges of $L_{1}(G)$ is $\Theta(nr^{d})$.
Proof.
The upper bound on the number of edges follows trivially, as a.a.s. the total number of edges in $G$ is $\Theta(nr^{d})$. For the lower bound, first sprinkle a Poisson point process of intensity $1-\xi$ where $\xi>0$ is small enough so that the resulting graph has a giant component of size $\Theta(n)$ a.a.s. Now, each vertex of this giant has in expectation $\Theta(r^{d})$ neighbors among vertices of the second PPP of intensity $\xi$, and so the total number of edges emanating from the giant component of the first PPP is $\Theta(nr^{d})$. By a standard second moment method (we may assume $r=O(\log n$)), concentration follows. ∎
We will make use of the following proposition that relates the Euclidean distance of two vertices of $L_{1}(G)$ to the graph distance. By Theorem 3 and Remark 5 of [21], we have:
Theorem 2.11.
Let $d\geq 2$, let $G\in\mathcal{G}(n,r,d)$ and let $r_{g}<r$. A.a.s. the following is true: there exist positive constants $C=C(r,d)$ and $C^{\prime}=C^{\prime}(r,d)$, such that for any two vertices $x,y\in L_{1}(G)$ with $d_{E}(x,y)\geq C\frac{\log n}{r^{d-1}}$, we have $d_{G}(x,y)\leq C^{\prime}d_{E}(x,y)/r$.
We get the following corollary:
Corollary 2.12.
Let $d\geq 2$, let $G\in\mathcal{G}(n,r,d)$ and let $r_{g}<r=O(r_{c})$. Let $V^{\prime}\subseteq L_{1}(G)$ be such that for any vertex $x\in L_{1}(G)$ we have that $d_{E}(x,V^{\prime})<C\log n/r^{d-1}$ for the same constant $C$ as in Theorem 2.11. Then, a.a.s., there exists a constant $c^{\prime}>0$ such that $d_{G}(x,V^{\prime})\leq c^{\prime}\log n/r^{d}$ for any vertex $x\in L_{1}(G)$.
Proof.
On the one hand, for every $x\in L_{1}(G)$ there exists $y\in V^{\prime}$ with $d_{E}(x,y)<C\log n/r^{d-1}$, but on the other hand, since $x\in L_{1}(G)$, a.a.s. there exists $v\in L_{1}(G)$ with $d_{E}(x,v)\geq n^{\varepsilon}$ for some $\varepsilon>0$. Thus, there exists $z\in L_{1}(G)$ such that $d_{E}(x,z)\geq C\log n/r^{d-1}$ and $d_{E}(x,z)=O(\log n/r^{d-1})$: indeed, consider a shortest path $x=v_{0},v_{1},\ldots,v_{k}=v$ for some $k\in\mathbb{N}$: since the Euclidean distance between two consecutive vertices is at most $r$, there exist vertices $v_{i}$ and $v_{i+1}$ on this path such that $d_{E}(x,v_{i})<C\log n/r^{d-1}$ and $d_{E}(x,v_{i+1})\geq C\log n/r^{d-1}$, but at the same time $d_{E}(x,v_{i+1})\leq C\log n/r^{d-1}+r=O(C\log n/r^{d-1})$, where the equality holds from our assumption $r=O(r_{c})$. Therefore, by Theorem 2.11, $d_{G}(x,v_{i+1})\leq c^{\prime}\log n/r^{d}$ for some $c^{\prime}>0$ and since $v_{i+1}\in V^{\prime}$, the statement follows. ∎
Next, we also use the following result (see for example [35], Theorem 10.18):
Theorem 2.13.
Let $d\geq 2$, let $G\in\mathcal{G}(n,r,d)$ and let $r>r_{g}$. A.a.s., there exists $c>0$ such that the second largest connected component of $G$ is of size at most $c\log^{2}n$.
2.3. Bernoulli percolation
We also need results from Bernoulli percolation. Consider an i.i.d. family of independent Bernoulli variables with mean $p$, indexed by the edges of $\mathbb{Z}^{d}$ (typically called bonds). We say that the bond $e$ is open if its associated variable is equal to 1. For a given realization, we generate a subgraph of $\mathbb{Z}^{d}$ with edges given by the open bonds, this random graph model is known as Bernoulli bond percolation. For $d\geq 2$, it is known that this model undergoes a non-trivial phase transition: there exists $10<p_{c}^{bond}<$ 1 such that if $p>p_{c}^{bond}$ then a.s. there is a unique (maximal) infinite connected component $\mathcal{C}$ (see for example [9]), which we call the infinite cluster. If one considers variables indexed by the vertices of $\mathbb{Z}^{d}$ instead of bonds we have Bernoulli site percolation. We say that a vertex is open if its associated variable is 1. We say that a path of vertices is open if all its vertices are open. As in bond percolation, for $d\geq 2$ there exists $10<p_{c}^{site}<$ 1 such that if $p>p_{c}^{site}$ then there is a.s. a unique infinite cluster.
We denote by $\mathbb{P}^{site}_{p}$ the law of Bernoulli site percolation with parameter $p$, and by $\mathbb{P}^{bond}_{p}$ the law of Bernoulli bond percolation with parameter $p$, respectively. We denote by $\mathcal{C}^{bond}(n)$ the largest connected component of Bernoulli bond percolation inside the box $\left[-n^{1/d}/2,n^{1/d}/2\right]^{d}$ if it exists. Equipping $\mathcal{C}^{bond}(n)$ with the network structure induced by the simple random walk, we have the following sharp estimate for the effective resistance of this network:
Theorem 2.14 (Theorem 1.1 of [1]).
Let $d\geq 2$. For bond percolation on $\mathbb{Z}^{d}$ and $p\in(p_{c}^{bond},1)$ there are constants $c_{1},c_{2}>0$ such that a.a.s., it is true that
$c_{1}\log n\leq\max_{x,y\in L_{1}^{bond}}\mathcal{R}\left(x\longleftrightarrow y,\mathcal{C}^{bond}(n)\right)\leq c_{2}\log n.$ (2.10)
We also make use of the following percolation estimations of crossings in boxes for 2-dimensional site percolation. Given $[0,m]\times[0,n]\subseteq\mathbb{Z}^{2}$, a horizontal crossing is an open and self-avoiding path entirely contained within the box such that the first vertex of the path has its first coordinate equal to 00 and the final vertex has its first coordinate equal to $m$ (path going from left to right). We define a vertical crossing in a similar way. A long crossing of a rectangle $[a,b]\times[c,d]$ is an open self-avoiding path along the longest dimension of the rectangle (ties broken arbitrarily). We will use the following result:
Lemma 2.15 (Theorem 11.1 of [27]).
Consider site percolation on $\mathbb{Z}^{2}$ with parameter $p>p_{c}^{site}$. There exist positive constants $c_{1},c_{2},c_{3}$ such that
$\begin{split}\mathbb{P}^{site}_{p}\big{(}&\text{There exist at least $c_{1}n$ disjoint horizontal crossings}\\ &\text{ in $([0,m]\times[0,n])\cap\mathbb{Z}^{2}$}\big{)}\geq 1-c_{2}me^{-c_{3}n}.\end{split}$ (2.11)
Finally, we need a result on the chemical distance for Bernoulli site percolation (see Lemma 2.3 of  [1]):
Theorem 2.16.
Let $d\geq 2$ and $p>p_{c}^{site}$. There exist constants $\kappa_{0},c$ such that a.a.s. the following holds: for any $x,y\in\mathcal{C}^{site}(n)$, and any $\kappa>\kappa_{0}$, if $d_{E}(x,y)\leq\kappa\log n$, then $d_{G}(x,y)\leq c\kappa\log n.$
3. Renormalization scheme
In this section we explain the renormalization scheme that we use in order to obtain the upper bound of Theorem 1.1. The objective of the renormalization used by Abe in  [1] is similar, here it is used to prove that there is a backbone of crossings of $L_{1}(G)$. We encounter several technical challenges proper of random geometric graphs that we have to consider. Fix throughout this section $M>0$ large enough. For each $\mathrm{i}\in\mathbb{Z}^{d}$ we define the cube
$\mathrm{C_{i}}=Mr\mathrm{i}+[-Mr/2,Mr/2]^{d}.$ (3.1)
That is, $\mathrm{C_{i}}$ is an axis-parallel cube with center $Mr\mathrm{i}$ and side length $Mr$. We say it has label $\mathrm{i}$. We say that two cubes $\mathrm{C_{i}}$ and $\mathrm{C_{j}}$ are neighbors if $\mathrm{i}\sim\mathrm{j}$ (on the renormalized lattice resulting from tessellation of $\mathbb{Z}^{d}$ into cubes of sidelength $Mr$). Denote by $\mathcal{K}_{n}$ the set of cubes that are totally contained within $\Lambda_{n}=\left[-\frac{n^{1/d}}{2},\frac{n^{1/d}}{2}\right]^{d}$. The choice of $M$ depends on the section: in Section 4 the choice of $M$ depends on another renormalization scheme that is introduced there. In the other sections, we may assume that by an appropriate choice $M$, all cubes $\mathrm{C_{i}}$ are either completely contained in $\Lambda_{n}$ or they share at most some faces with $\Lambda_{n}$.
Now, let $\mu=\mu(d)$ be the largest constant so that any two vertices in adjacent cubes (that is, sharing a $(d-1)-$dimensional face) of side length $r/\mu$ are connected by an edge (for example, in $d=2$, $\mu=\sqrt{5}$). We will consider a tessellation using such cubes $\mathrm{Q_{i}}$ of side length $r/\mu$ later on as well, for now we use it do define crossings. Let $U=[a_{1},a_{2}]\times[b_{1},b_{2}]\times U_{d-2}$ be a parallelepiped in $\mathbb{R}^{d}$, with $U_{d-2}$ a parallelepiped in $\mathbb{R}^{d-2}$. We say that the parallelepiped $U$ has a crossing component, if there exists a connected component (of the underlying $G\in\mathcal{G}(n,r,d)$) within the parallelepiped that contains vertices at Euclidean distance at most $\frac{r}{\mu}$ from each of the faces of the parallelepiped. Similarly as we did for $\mathbb{Z}^{d}$, we define also crossings: a horizontal crossing of $G\in\mathcal{G}(n,r,d)$ in $U$ is a self-avoiding path of $G$ such that the vertices of the path are totally contained in $U$, the first vertex of the path has first coordinate in the interval $(a_{1},a_{1}+r/\mu)$, and the last vertex of the path has first coordinate in the interval $(a_{2}-r/\mu,a_{2})$. A vertical crossing is defined in the same way, exchanging the first coordinate by the second coordinate and $a_{1},a_{2}$ by $b_{1},b_{2}$, and as before, we use also the definition of a long crossing for a crossing in the longer dimension. We remark that while a horizontal (or vertical) crossing is a self-avoiding path, a crossing component is in general not a path (but it contains such a path).
We then define the indicator event $U_{\mathrm{i}}$ related to cube $\mathrm{C_{i}}$ to be true iff the following two conditions hold:
  1. (a)
    $\mathrm{C_{i}}\cup\mathrm{C_{j}}$ contains a crossing component for each $\mathrm{j}\sim\mathrm{i}$
  2. (b)
    $\mathrm{C_{i}}$ has only one connected component with diameter larger than $Mr/5$.
Ci
Figure 1. The central cube $C_{i}$ and its surrounding paths. The colored lines represent vertex paths of $G\in\mathcal{G}(n,r,d)$.
$\Lambda_{n}$$\mathcal{K}_{n}^{2}$
Figure 2. A $2$-dimensional slice $\mathcal{K}_{n}^{2}$, with a red $\alpha$-logarithmic strip.
By a straight-forward modification of a result of Penrose and Pisztora (Theorem 2 of [37]) (in the original definition of Penrose and Pisztora, the crossing component is at distance at most $r$ from all the faces of the parallelepiped, but the proof works in exactly the same way for $r/\mu$) we may state the following lemma:
Lemma 3.1.
Let $d\geq 2$, let $M$ be a large enough constant, let $G\in\mathcal{G}(n,r,d)$ and let $r\geq(1+\varepsilon)r_{g}$ for arbitrarily small $\varepsilon>0$. Fix $\mathrm{i}\in\mathbb{Z}^{d}$. If $\mathrm{i}\sim\mathrm{j}$, there exists a constant $\gamma>0$ such that
$\begin{split}\mathbb{P}&(\mathrm{C_{i}}\cup\mathrm{C_{j}}\text{ has a crossingcomponent})\geq 1-e^{-\gamma Mr^{d}},\\ \mathbb{P}&(\mathrm{C_{i}}\text{ has only one connected component with diameter larger than $Mr/5$})\geq 1-e^{-\gamma Mr^{d}}.\end{split}$ (3.2)
Therefore, for $M$ large enough there is a constant $\gamma^{\prime}>0$ such that for any $\mathrm{i}\in\mathbb{Z}^{d}$ we have $\mathbb{P}(U_{\mathrm{i}}=1)\geq 1-e^{-\gamma^{\prime}Mr^{d}}$. If $U_{\mathrm{i}}=1$ we say that the cube $\mathrm{C_{i}}$ is good (see Figure 2). Note that the indicator events $U_{\mathrm{i}}$ and $U_{\mathrm{j}}$ are independent if $\mathrm{i},\mathrm{j}$ are at graph distance (in the renormalized lattice graph) of at least 2. By Theorem 0.0 of [31], for $M$ large enough, $(U_{\mathrm{i}})_{\mathrm{i}\in\mathbb{Z}^{d}}$ stochastically dominates a family $(Y_{\mathrm{i}})_{\mathrm{i}\in\mathbb{Z}^{d}}$ of i.i.d. Bernoulli variables with parameter $p$, where $p\to 1$ as $M\to\infty$. In particular, seeing the good cubes as open sites of Bernoulli site percolation, the set of good cubes thus has a giant component. Moreover, if a cube $\mathrm{C_{i}}$ is good and $\mathrm{C_{i}}$ is in the giant component of good cubes, then any vertex $x\in\mathrm{C_{i}}$ belonging to the largest component of $G$ restricted to $\mathrm{C_{i}}$ belongs to $L_{1}(G)$: indeed, by repeated application of Lemma 3.1, the giant component of the renormalized lattice contains, since the component with large diameter is unique, a.a.s. a long crossing of order $\Omega(n^{1/d}/r)\geq n^{\varepsilon}$ for some $\varepsilon>0$ (we may assume $r=O(\log n)$ throughout the paper), which by Theorem 2.13 a.a.s. is the giant component of $G$.
A 2-dimensional slice of $\mathcal{K}_{n}$ is a subset of cubes with labels $\mathrm{i}$ such that all but $2$ coordinates are fixed (see Figure 2 for an example). We denote such a slice by $\mathcal{K}_{n}^{2}$. Since very $2$-dimensional slice is isomorphic to $\mathbb{Z}^{2}$ restricted to a box of a certain side length, by abuse of notation we can thus also define both vertical and horizontal orientations for each slice. We remark also, that, as before, the good cubes of $\mathcal{K}^{2}_{n}$ dominate supercritical 2-dimensional site percolation restricted to a large box.
Now, divide a 2-dimensional slice $\mathcal{K}_{n}^{2}$ into vertical disjoint strips of cubes $\mathrm{C_{i}}$ of width $\lceil\alpha\log n\rceil$. The total Euclidean width of the strip is $\lceil\alpha\log n\rceil Mr$ (as each cube has side length $Mr$), whereas its Euclidean diameter (that is, the largest Euclidean distance between any two points in such a strip) is $\Theta(n^{1/d})$, see Figure  2 (the strips that are at the boundary of the slice have a smaller width). We call them $\alpha$-logarithmic vertical strips (or simply strips, in case the orientation is not important), and analogously we define $\alpha$-logarithmic horizontal strips.
We prove the following result regarding long crossings of good cubes inside of $\Lambda_{n}$:
Lemma 3.2.
Let $d\geq 2$ and $\varepsilon>0$, let $G\in\mathcal{G}(n,r,d)$ and assume $(1+\varepsilon)r_{g}\leq r=O(r_{c})$ and assume $M$ large enough. Let $\alpha$ be a large enough positive constant. Then, a.a.s. there exists $c>0$ such that each $\alpha$-logarithmic strip that is not in the boundary of $\Lambda_{n}$ of every 2-dimensional slice of $\Lambda_{n}$ has at least $c\log n$ disjoint long crossings of good cubes.
Proof.
As noted after Lemma  3.1, for $M$ large enough, the good cubes of the tessellation stochastically dominate a super-critical Bernoulli site percolation process. Observe that the number of slices of $\Lambda_{n}$ is $O(n^{1/d})$, and also the number of $\alpha$-logarithmic strips inside a slice is $O(n^{1/d})$. Hence, by Lemma 2.15 applied to every $\alpha$-logarithmic strip, together with a union bound, the probability that at least one strip does not have $c\log n$ long crossings of good cubes is at most $O(n^{1/d})O(n^{1/d}c_{2}n^{1/d}e^{-c_{3}\alpha\log n})=o(1)$, where we assumed $\alpha$ large enough. ∎
4. Lower bound on the cover time
Recall that $\mu=\mu(d)$ is the largest constant so that any two vertices in adjacent cubes (that is, two cubes sharing a $d-1$-dimensional face) of sidelength $r/\mu$ are connected by an edge. We are going to use the tessellation into cubes $\mathrm{C_{i}}$ as described in Section 3 of side length $Mr$ and introduce moreover a new finer tessellation of $\Lambda_{n}$ into cubes $\mathrm{Q_{i}}$, of the already mentioned sidelength $\frac{r}{\mu}$. The cube with index $\mathrm{i}\in\mathbb{Z}^{d}$ is given by:
$\mathrm{Q_{i}}=\mathrm{i}\frac{r}{\mu}+\left[0,\frac{r}{\mu}\right]^{d}$ (4.1)
Let $\phi(d)=\phi$ be the smallest integer such that any pair of vertices in $G\in\mathcal{G}(n,r,d)$, contained in cubes $\mathrm{Q_{i}}$ and $\mathrm{Q_{j}}$, is not connected by an edge if $|\mathrm{i}-\mathrm{j}|_{1}>\phi$. For example, $\phi(2)=3$. Let $x\in\Lambda_{n}$ be a vertex of $G\in\mathcal{G}(n,r,d)$ with $x\in\mathrm{Q_{j}}$. We say that $x$ is $m$-special if the following two conditions hold:
  • Each of the cubes $\mathrm{Q_{k}}$, with $\mathrm{k}=\mathrm{j}+e_{1}s$ and $1\leq s\leq m$, has exactly one point of the PPP. We call the set of cubes $\{\mathrm{Q_{k}}\}$ by slight abuse of notation $m$-pending path with root at $x$ (note that the graph induced by a pending path is not necessarily a path, since vertices in two cubes at distance $2$ in the cube metric might also be connected by an edge).
  • All cubes at distance less or equal to $\phi$ to the pending path in the |.$|.|_{1}$ norm (where the cubes are seen as points of $\mathbb{Z}^{d}$) and to the right of $\mathrm{Q_{j}}$ are empty (see Figure 3 for an example).
Tessellate $\Lambda_{n}$ into cubes of side length $Mr$ as in Section 3, such that $Mr$ is large enough and also an integer multiple of $r/\mu$ (in this way cubes $\mathrm{Q}$ are either totally contained in cubes $\mathrm{C}$ or they share faces or they are disjoint). Note that such a choice of $M$ only depends on the dimension. We only consider cubes of sidelength $Mr$ fully contained in $\Lambda_{n}$; it might happen that close to the boundary some part of $\Lambda_{n}$ remains uncovered by the tessellation, but this causes no problem with the proofs of this section.
The following lemma tells us that for $c_{1}>0$ small enough, there are relatively many $\left\lfloor c_{1}\frac{\log n}{r^{d}}\right\rfloor$-pending paths, whose root is attached to the giant of $G\in\mathcal{G}(n,r,d)$.
Lemma 4.1.
Let $d\geq 2$, $\varepsilon>0$, and let $G\in\mathcal{G}(n,r,d)$. There exist small constants $c_{1},\xi,\gamma$ ($\xi$ depends only on $d,c_{1},\gamma$), such that if $(1+\varepsilon)r_{g}\leq r<\gamma r_{c}$, the event
$\left\{\left|x\in L_{1}(G):x\text{ is }\left(\left\lfloor c_{1}\frac{\log n}{r^{d}}\right\rfloor\right)-\mbox{special}\right|\geq n^{\xi}\right\}$ (4.2)
occurs a.a.s.
Proof.
In order to make sure that a $\left\lfloor c_{1}\frac{\log n}{r^{d}}\right\rfloor$-pending path is completely contained in $\Lambda_{n}$, we would like its root to be inside $\Lambda_{m_{n}}$, where
$m_{n}=n^{1/d}-\frac{r}{\mu}\left\lfloor c_{1}\frac{\log n}{r^{d}}\right\rfloor$ (4.3)
Observe that by Lemma 2.13, a.a.s. all vertices that are in a horizontal crossing of $\Lambda_{m_{n}}$ belong to $L_{1}(G)$, as a crossing in $\Lambda_{m_{n}}$ is a connected component of size $\Omega(n^{1/d})$. Since we are interested in a result that holds a.a.s., we may condition on this.
Let $\mathcal{K}_{m_{n}}^{2}$ be the 2-dimensional slice where the first two coordinates are not fixed and all others are equal to 0 (see Figure 2). By Lemma 3.2, a.a.s., there are at least $c_{2}m_{n}$ horizontal disjoint crossings of good cubes inside $\mathcal{K}_{m_{n}}^{2}$ for some $c_{2}>0$, and as before, we may condition on this as well. Each horizontal crossing of good cubes contains a crossing of $G$ inside $\Lambda_{m_{n}}$ (recall that the right endpoints of the crossing in $G$ are at distance less than $r/\mu$ from the right boundary of the slice $\mathcal{K}_{m_{n}}^{2}$).
Let $x\in\mathrm{Q_{i}}$ and $y\in\mathrm{Q_{j}}$ be vertices of $G\in\mathcal{G}(n,r,d)$, with $\mathrm{i}_{1}=\mathrm{j}_{1}$ (that is, with first coordinates being equal). Note that if $|\mathrm{i}-\mathrm{j}|>2\phi$ (see the definition of $\phi$ above) then the events of the vertices $x,y$ being $\left(\left\lfloor c_{1}\frac{\log n}{r^{d}}\right\rfloor\right)$-special are independent. Also, note that the configuration inside $\mathcal{K}_{m_{n}}$ is independent of what happens in $\Lambda_{n}\setminus\mathcal{K}_{m_{n}}$. In particular, the event that a vertex $x$ that is inside of $\mathcal{K}^{2}_{m_{n}}$, is at distance less than $r/\mu$ from the right boundary of $\mathcal{K}^{2}_{m_{n}}$ and belongs to $L_{1}(G)$ does not have an effect on the probability to have a $\left\lfloor c_{1}\frac{\log n}{r^{d}}\right\rfloor$-pending path in $\Lambda_{n}\setminus\mathcal{K}^{2}_{m_{n}}$ attached to it. Using the fact that $\lfloor c_{1}\log n/r^{d}\rfloor\geq 1$, the probability for the latter event to happen is equal to
$e^{-(r/\mu)^{d}}(r/\mu)^{d})^{\left\lfloor c_{1}\frac{\log n}{r^{d}}\right\rfloor}(e^{-(/\mu)^{d}})^{c^{\prime}\left\lfloor c_{1}\frac{\log n}{r^{d}}\right\rfloor}(e^{-(r/\mu)^{d}})^{c^{\prime\prime}}\geq n^{-\rho},$ (4.4)
where $c^{\prime},c^{\prime\prime}$ are constants that depend only on the dimension and $\rho$ is a constant that depends on $c_{1},d,\varepsilon,\gamma$. Observe that we can choose $c_{1}$ small enough such that $\rho<1/d$, and then we can choose $\gamma$ small enough such that $c_{1}\frac{\log n}{r^{d}}$ is always greater than 1. Hence, using the fact that a.a.s. there are at least $c_{3}n^{1/d}$ right endpoints of horizontal crossings of $L_{1}(G)$ in $\Lambda_{m_{n}}$ (taken from long crossings of good cubes, and that are at distance less than $r/\mu$ of the right boundary of the slice), and that at least a constant fraction of all of them give rise to independent events of being $\lfloor c_{1}\log n/r^{d}\rfloor$-special, we have
({|$\begin{split}\mathbb{P}\left(\{|\{x\in L_{1}(G):x\text{ is }(\lfloor c_{1}\log n/r^{d}\rfloor)-\mbox{special}\}|<n^{\xi}\}\right)\leq\mathbb{P}\left(Bin\left(c_{4}n^{1/d},n^{-\rho}\right)<n^{\xi}\right).\end{split}$ (4.5)
By choosing $\xi=(1/d-\rho)/2$ and applying Chernoff bounds, the theorem follows. ∎
Theorem 4.2.
Let $\varepsilon$ be an arbitrarily small positive constant. Let $(1+\varepsilon)r_{g}\leq r\leq(1-\varepsilon)r_{c}$. Then a.a.s.,
$\tau_{cov}(L_{1}(G))\geq\delta n\log^{2}n,$ (4.6)
where $\delta$ is a constant that only depends on $\varepsilon$.
Proof.
Case 1: $r<\gamma^{\prime}r_{c}$, where $\gamma^{\prime}\leq\gamma$ ($\gamma$ is the same constant from Lemma 4.1) is a small enough constant to be fixed later. Define as before
$m_{n}=n^{1/d}-\frac{r}{\mu}\left\lfloor\frac{c_{1}\log n}{r^{d}}\right\rfloor.$ (4.7)
Let $\mathcal{P}$ be the set of $\left\lfloor\frac{c_{1}\log n}{r^{d}}\right\rfloor$-special vertices of $\Lambda_{m_{n}}$ that are contained within $L_{1}(G)$ (restricted to $\Lambda_{m_{n}})$. Denote by $\mathcal{F}$ the set of vertices of $L_{1}(G)$ that are in the rightmost cube of a pending path with root at a vertex in $\mathcal{P}$ (see Figure 3):
=$\mathcal{F}=\{y\in L_{1}(G):y\in\mathrm{C_{j}},\,\mathrm{j}=\mathrm{i}+\lfloor c_{1}\log n/r^{d}\rfloor,x\in\mathrm{Q_{i}}\text{ and }x\in\mathcal{P}\}$ (4.8)
Note that $|\mathcal{P}|=|\mathcal{F}|$. By Lemma 4.1, $|\mathcal{F}|\geq n^{\xi}$ a.a.s., and by construction, $\mathcal{F}\subseteq L_{1}(G)$. Now, we give a lower bound for the effective resistance between any pair of vertices of $\mathcal{F}$. Take $x,y\in\mathcal{F}$, and let us construct edge-disjoint edge-cutsets separating $x$ from $y$. Let $x\in\mathrm{Q_{i}}$. Let $\sigma(d)=\sigma$ be the minimum integer such that given $\sigma$ consecutive cubes of a pending path, any path of $G$ that connects any two vertices in different components of the pending path when taking out vertices in these $\sigma$ cubes must have at least one edge with both endpoints inside the $\sigma$ many consecutive cubes. For $11\leq k\leq\left\lfloor\frac{1}{\sigma}\left\lfloor\frac{c_{1}\log n}{r^{d}}\right\rfloor\right\rfloor-$ 1 (the $-1$ is to avoid boundary cases that require separate calculations), let $\Pi_{k}$ be the set of edges of $L_{1}(G)$ with both endpoints are inside the cube $\mathrm{Q_{j}}$ with $\mathrm{j}=\mathrm{i}-e_{1}s$ and $\sigma(k-1)\leq s<\sigma k$. Choose $\gamma^{\prime}$ small enough such that $\left\lfloor\frac{1}{\sigma}\left\lfloor\frac{c_{1}\log n}{r^{d}}\right\rfloor\right\rfloor$ is greater or equal to 2. Note that $|\Pi_{k}|\leq\binom{\sigma}{2}$, as there is only one vertex per cube $\mathrm{Q_{j}}$. Moreover, observe that $\Pi:=\{\Pi_{k}\}_{k}$ is a family of disjoint edge-cutsets between $x,y$, as any path between both vertices has to contain an edge belonging to $\Pi_{k}$ (by definition of $\sigma$), for any $k$. By Lemma 2.5 we then have
$\mathcal{R}(x\longleftrightarrow y,L_{1}(G))\geq\sum_{k=1}^{\left\lfloor\frac{1}{\sigma}\left\lfloor\frac{c_{1}\log n}{r^{d}}\right\rfloor\right\rfloor-1}\frac{1}{\binom{\sigma}{2}}\geq\delta_{1}\frac{\log n}{r^{d}},$ (4.9)
for some $\delta_{1}>0$. Using Lemma  2.10 together with the formula of Theorem 2.7 we obtain the desired result for this case.
Case 2: $\gamma^{\prime}r_{c}\leq r\leq(1-\varepsilon)r_{c}$. In this case we will use a different strategy: we sill show that a.a.s. the number of vertices of $L_{1}$ with degree 1 is $\Omega(n^{\xi^{\prime}})$ for some $\xi^{\prime}>0$. First, note that the expected number of vertices $v\in G$ having degree exactly $1$, and having the only neighbor at radial distance at least $r/2$ is at least $(1+o(1))n^{1-\xi^{\prime}}$ for some $10<\xi^{\prime}<$ 1: indeed, for each vertex, by our assumption on $r$, the probability to have such only neighbor is $e^{-c_{1}(r/2)^{d}}c_{2}r^{d}e^{-c_{3}r^{d}}=n^{-\xi^{\prime}}$ for some dimension-dependent $c_{1},c_{2},c_{3}>0$ and $10<\xi^{\prime}<$ 1, and the statement follows since a.a.s. there are $n(1+o(1))$ vertices. Moreover, by a standard second moment method, there are a.a.s. $(1+o(1))n^{1-\xi^{\prime}}$ such vertices. The same holds when restricting to vertices at distance at least say $r^{2}$ (the value of $r^{2}$ is arbitrary, only to make sure that paths below remain inside $\Lambda_{n}$) from all boundaries of $\Lambda_{n}$, and we will assume this in the sequel.
We need to show that for each vertex there is at least a constant chance to be part of $L_{1}$. Denote first by $C_{d}$ be the smallest dimension-dependent integer such that any cube of a given tessellation has at most $C_{d}$ adjacent cubes in the tessellation (that is, cubes sharing a lower-dimensional face). We will now show that every component different from $L_{1}$, by our choice of $r$, a.a.s. has Euclidean diameter at most $O(r)$ in $G$, or equivalently, Euclidean diameter $O(1)$ on the renormalized lattice: indeed, It is a standard topological fact that in the tessellation of side length $r/\mu$, such component has to be separated by a diagonally connected set of cells from $L_{1}$, all being empty. The expected number of such components with more than $L$ cells is at most $n(C_{d}e^{-V_{d}(r/\mu)^{d}})^{L}$. Plugging in our assumption on $r$ and sufficiently large $L=L(d)$, we see that this expectation is $o(1)$, and hence, by Markov’s inequality there is a.a.s. no such set. Therefore, a.a.s. every other component has Euclidean diameter at most $Lr$, and we condition on this as well.
Next, having explored the neighborhood of one vertex $v=v_{0}$ and having found exactly one vertex $v_{1}$ at Euclidean distance at least $r/2$, we see that the ball of radius $r$ around $v_{1}$ that was not yet explored is of area at least $c^{\prime\prime}r^{d}$ for some $c^{\prime\prime}>0$. Assume w.l.o.g. that the $x$-coordinate of $v_{1}$ is at least $r/2$ larger than the one of $v_{0}$. Conditional on this, the probability that $v_{1}$ has at least one neighbor with $x$-coordinate at least $r/2$ larger than the one of $v_{1}$ is again $1-n^{-\xi^{\prime\prime}}$ for some $\xi^{\prime\prime}>0$, since this area had not been explored before and the unexplored area is $\Theta(\log n)$. Conditional under having found such a vertex, say $v_{2}$, we proceed iteratively: having found $v_{i}$ in this way, we look again for at least one neighbor $v_{i+1}$ with $x-$coordinate at least $r/2$ larger than $v_{i}$. By the same argument, independently of the found vertices, the probability to find such a vertex is at least $1-n^{-\xi^{\prime\prime}}$. Hence, with $L$ from before, by a union bound over all steps, with probability at least $1-L^{2}n^{-\xi^{\prime\prime}}=1-o(1)$ we find at least $L^{2}$ consecutive vertices in the same component as $v_{0}$. We advance at least $L^{2}r/2>Lr$ in the $x$-direction, and conditional under the statements before, $v_{0}$ must therefore belong to $L_{1}$. Hence, with probability $1-o(1)$, every vertex of degree $1$, whose neighbor is at distance at least $r/2$, belongs to $L_{1}$. Thus, the expected number of vertices of $L_{1}$ having degree $1$ is at least $(1-o(1))n^{1-\xi^{\prime}}$. Since for every vertex the $L^{2}$-neighborhood a.a.s. contains a polylogarithmic number of vertices, another second moment method shows that this also holds a.a.s.
To conclude, let $V^{\prime}$ then be the set of vertices of $L_{1}$ of degree exactly $1$. By the above, a.a.s. $|V|=\Omega(n^{1-\xi^{\prime}})$ for some $10<\xi^{\prime}<$ 1. For any different $x,y\in V^{\prime}$, taking out the only edge incident to say $x$, we obtain an edge-cutset of size $1$ between $x$ and $y$, and thus, by Lemma 2.5,
$\mathcal{R}(x\longleftrightarrow y,L_{1}(G))\geq 1.$
Since this holds for any different $x,y\in V^{\prime}$, using Lemma 2.10, by Theorem 2.7, a.a.s. there exist some $\delta_{2},\delta_{3}>0$, such that
$\tau_{cov}\geq\delta_{2}nr^{d}\log n\geq\delta_{3}n\log^{2}n,$
and this case is finished as well, and by choosing $\delta=\min\{\delta_{1},\delta_{3}\}$ the proof of the theorem is finished. ∎
[…]
Figure 3. A pending path in 2 dimensions with root at $x$ ($x$ is thus $m$-special for some $m$); all cubes surrounding the path are empty; $y$ is the vertex associated to $x$ that belongs to $\mathcal{R}$. Observe that a vertex in the path could be connected by an edge to more than 2 vertices (the maximum being 5 vertices).
5. Upper bound on the cover time for fixed radius
To prove the upper bound, we first introduce a structure of random paths between good cubes of a 2-dimensional slice. This construction was first introduced by Abe  [1]. Our first step will then be to show that for any vertices $x,y$ belonging to corresponding good cubes $\mathrm{C_{i}}$ and $\mathrm{C_{j}}$ as well as to the largest components of these cubes, and therefore in particular belonging to $L_{1}(G)$, we can suitably bound from above the resistance between $x$ and $y$. We thus focus in the next subsection on good cubes only. Throughout this section, we consider the tessellation of cubes $\mathrm{C_{i}}$ introduced in Section 3. In this section we will allow $M$ to change slightly with $n$ so that a cube $\mathrm{C_{i}}$ is either totally contained within $\Lambda_{n}$ or shares at most some faces with $\Lambda_{n}$, and so $\Lambda_{n}$ is completely covered by the tessellation. Given an integer $M$, we can find for each $n$, the closest value to it that satisfies this is at most $O(1/n)$ away, which will not play any role below.
5.1. Random path of good cubes
We assume $r$ fixed for the rest of this subsection.
Given two self-avoiding paths $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ of cubes (with direction given by the ordering of the cubes in the path) and $\mathrm{C}\in\mathcal{L}_{1}\cap\mathcal{L}_{2}$, the concatenation of both paths at $\mathrm{C}$ is the path obtained from starting at $\mathcal{L}_{1}$ until hitting $\mathrm{C}$ and following $\mathcal{L}_{2}$ after that. Note that if $\mathrm{C}$ is the last cube of intersection between the two paths (seen from $\mathcal{L}_{2}$) the concatenation is self-avoiding.
In the same way we define the concatenation of two self-avoiding paths $\mathcal{J}_{1}$ and $\mathcal{J}_{2}$ of $G\in\mathcal{G}(n,r,d)$, given at a vertex $x$ that belongs to both paths, as the path that results from starting at $\mathcal{J}_{1}$ until hitting $x$ and after that continuing with $\mathcal{J}_{2}$.
Recall that by Lemma  3.2, for $\alpha$ and $M$ large enough, a.a.s. there are at least $\lfloor c\log n\rfloor$ long crossings of good cubes in each $\alpha$-logarithmic strip (that is not in the boundary) of every 2-dimensional slice of $\Lambda_{n}$. In what follows we condition on this event.
Without loss of generality suppose in the following that $\mathcal{K}^{2}_{n}$ is the 2-dimensional slice made out of cubes $\mathrm{C_{i}}$ with only the first two coordinates not equal to 0. Divide $\mathcal{K}_{n}^{2}$ into disjoint strips of $\alpha$-logarithmic size (see Figure  2). Recall that for a cube $\mathrm{C_{j}}$, we denote by $\mathrm{j}_{1},\ldots,\mathrm{j}_{d}$ its $d$ coordinates. For $s\in\mathbb{Z}$, denote the $s$-th $\alpha$-logarithmic horizontal strip by
$\mathrm{H}_{s}=\{\mathrm{C_{j}}\in\mathcal{K}_{n}^{2}:\lceil\alpha\log n\rceil-s\leq\mathrm{j}_{2}<s+\lceil\alpha\log n\rceil\},$ (5.1)
and similarly denote the $s$-th $\alpha$-logarithmic vertical strip by
$\mathrm{V}_{s}=\{\mathrm{C_{j}}\in\mathcal{K}_{n}^{2}:\lceil\alpha\log n\rceil-s\leq\mathrm{j}_{1}<s+\lceil\alpha\log n\rceil\}$ (5.2)
Recall also that strips at the boundary of the slice might have a smaller width. Select $\lfloor c\log n\rfloor$ disjoint long crossings of good cubes for each strip $\mathrm{H}_{s},\mathrm{V}_{s}$ that is not on the boundary. We enumerate the vertical crossings of good cubes of the vertical strip $\mathrm{V}_{s}$, counting from right to left and denote them by $\mathrm{V}_{s}^{\ell}$ for $1\leq\ell\leq\lfloor c\log n\rfloor$ (this way of enumerating the vertical crossings is going to be important later). Similarly, we enumerate the horizontal crossings of $\mathrm{H}_{s}$, this time counting from bottom to top and denoting them by $\mathrm{H}_{s}^{\ell}$, for $1\leq\ell\leq\lfloor c\log n\rfloor$.
For $r,s\in\mathbb{Z}$, denote by $\mathrm{T}_{s,r}$ the smallest parallelepiped that contains all cubes $\mathrm{C_{i}}$ from $\mathrm{V}_{s}\cap\mathrm{H}_{r}$. Note that the parallelepiped $\mathrm{T}_{s,r}$ that is not at the boundary of the 2-dimensional slice has Euclidean side length $Mr\lceil\alpha\log n\rceil$ in the first 2 dimensions and $Mr$ in the others, as the side length of a cube $\mathrm{C_{i}}$ is $Mr$ (we consider in the construction that follows only parallelepipeds that are not at the boundary, this will be sufficient for our purposes). Also note that any vertical crossing of good cubes $\mathrm{V}^{\ell}_{s}$ of $\mathrm{V}_{s}$ and any horizontal crossing of good cubes $\mathrm{H}^{\ell}_{r}$ of $\mathrm{H}_{r}$ intersect in at least one good cube $\mathrm{C}$ that lies inside of $\mathrm{T}_{s,r}$, as the slice is a 2-dimensional object. Let now $\mathcal{L}$ be a self-avoiding path of parallelepipeds $\mathrm{T}$ between $\mathrm{T}_{s_{1},r_{1}}$ and $\mathrm{T}_{s_{2},r_{2}}$ (that is, a sequence of adjacent parallelepipeds of the same slice such that no parallelepiped appears more than once). We then construct an ordered set $(\mathrm{T}_{a_{1},b_{1}},\mathrm{T}_{a_{2},b_{2}},...,\mathrm{T}_{a_{k},b_{k}})$ of parallelepipeds that belong to $\mathcal{L}$ and that we use to identify the change of its direction in the following way: set $\mathrm{T}_{a_{1},b_{1}}=\mathrm{T}_{s_{1},r_{1}}$. Let $\mathrm{T}_{a_{l},b_{l}}\in\mathcal{L}$, and suppose without loss of generality that the next step of the path $\mathcal{L}$ is to the right. $\mathrm{T}_{a_{l+1},b_{l+1}}$ is then defined as the last parallelepiped of $\mathcal{L}$ that is visited by going only to the right of $\mathrm{T}_{a_{l},b_{l}}$. The construction is analogous if the path $\mathcal{L}$ after $\mathrm{T}_{a_{l},b_{l}}$ goes to the left, up or down, respectively. We finish at $\mathrm{T}_{a_{k},b_{k}}=\mathrm{T}_{s_{2},r_{2}}$. If $(a_{i+1},b_{i+1})-(a_{i},b_{i})=me_{1}$ for some $m$, we say that $\mathcal{L}$ has horizontal orientation at step $(a_{i},b_{i})$, and analogously, we say that it has vertical orientation if instead of $e_{1}$ we have $e_{2}$. See Figure  4 for an example of the construction.
[…]
Figure 4. Path $\mathcal{L}$ of parallelepipeds $\mathrm{T}$. The path changes direction when visiting the parallelepipeds $\mathrm{T}_{a_{i},b_{i}}$
Assume that $r_{1}=r_{2}$ for the parallelepipeds $\mathrm{T}_{s_{1},r_{1}}$ and $\mathrm{T}_{s_{2},r_{2}}$, and assume that $s_{2}>s_{1}$, so that $\mathrm{T}_{s_{2},r_{2}}$ is to the right of $\mathrm{T}_{s_{1},r_{1}}$. We will also throughout this subsection require $s_{2}-s_{1}$ to be odd, so that the parallelepipeds $\mathrm{T}_{s_{1},r_{1}}$ and $\mathrm{T}_{s_{2},r_{2}}$ are separated by an even number of parallelepipeds: we will later lift these restrictions. We now construct a self-avoiding random path $\mathcal{L}_{X}$ (based on the straight path $\mathcal{L}$) of parallelepipeds between $\mathrm{T}_{s_{1},r_{1}}$ and $\mathrm{T}_{s_{2},r_{2}}$: let $X$ be a continuous random variable chosen uniformly in $[0,\frac{s_{2}-s_{1}+1}{2}]$, and let $y_{X}\in\Lambda_{n}$ be the point that results from summing $(0,Mr\lceil\alpha\log n\rceil X,0,...,0)$ to the left bottom corner of $\mathrm{T}_{r_{1},\frac{s_{2}+s_{1}+1}{2}}$ (the $x$-coordinate of this point is the midpoint of $\mathrm{T}_{s_{1},r_{1}}$ and $\mathrm{T}_{s_{2},r_{2}}$, see Figure  5). Note that we assume here that $y_{X}$ is inside of $\Lambda_{n}$; if this were not the case we subtract $(0,Mr\lceil\alpha\log n\rceil,0,...,0)$ from the midpoint of $\mathrm{T}_{s_{1},r_{1}}$ and $\mathrm{T}_{s_{2},r_{2}}$, and the construction that follows is reflected vertically.
We denote by $\mathcal{L}_{X}$ the a.s. unique random path of parallelepipeds intersecting the segment between the bottom left corner of $\mathrm{T}_{s_{1},r_{1}}$ and the point $y_{X}$, and the segment between the bottom right corner of $\mathrm{T}_{s_{2},r_{2}}$ and $y_{X}$ (see Figure 6).
[…]
Figure 5. Position of $y_{X}$ relative to the parallelepipeds
[…]
Figure 6. The red cubes form the path $\mathcal{L}_{X}$
Let $\mathrm{C_{i}}\in\mathrm{T}_{s_{1},r_{1}}$ and $\mathrm{C_{j}}\in\mathrm{T}_{s_{2},r_{2}}$ be any good cubes such that both of them belong to some vertical crossings of the strips $\mathrm{V}_{s_{1}}$ and $\mathrm{V}_{s_{2}}$, respectively (see the leftmost part of Figure 7 for an example of such crossings for $\mathrm{T}_{s_{1},r_{1}}$). Either such crossings have length $O(\log n)$, or otherwise, by Theorem  2.16, a.a.s. we may select crossings with this property. We denote the selected vertical crossings by $\mathrm{V}(\mathrm{C_{i}})$ and $\mathrm{V}(\mathrm{C_{j}})$. From the vertical crossing of good cubes $\mathrm{V}(\mathrm{C_{i}})$ that contains $\mathrm{C_{i}}$ we then extract a bottom-to-top crossing of $\mathrm{T}_{s_{1},r_{1}}$ made from good cubes which we denote by $\mathrm{V}^{\prime}(\mathrm{C_{i}})$ (see yellow cubes in the middle picture of Figure 7). Similarly, we extract a bottom-to-top-crossing of $\mathrm{T}_{s_{2},r_{2}}$ made from good cubes, from the vertical crossing $\mathrm{V}(\mathrm{C_{j}})$ that contains $\mathrm{C_{j}}$, and we denote it by $\mathrm{V}{{}^{\prime}}(\mathrm{C_{j}})$. Given $1\leq\ell\leq\lfloor c\log n\rfloor$, we construct a self-avoiding path of good cubes between $\mathrm{C_{i}}$ and $\mathrm{C_{j}}$ that is entirely contained within $\mathcal{L}_{X}$ and has index $\ell$.
Since $\mathrm{V}{{}^{\prime}}(\mathrm{C_{i}})$ is self-avoiding, we may split it into two paths of good cubes with intersection only at $\mathrm{C_{i}}$. The paths in question are the ones originated from starting at $\mathrm{C_{i}}$ and going up or down following $\mathrm{V}{{}^{\prime}}(\mathrm{C_{i}})$, see the rightmost picture of Figure 7. Denote the paths by $\mathrm{V}_{1}^{\prime}(\mathrm{C_{i}})$ and $\mathrm{V}_{2}^{\prime}(\mathrm{C_{i}})$ and do the same for $\mathrm{V}^{\prime}(\mathrm{C_{j}})$. At least one of these paths of good cubes that start at $\mathrm{C_{i}}$ intersects $\mathrm{H}_{r_{1}}^{\ell}$ as the bottom-to-top crossing (of $\mathrm{T}_{s_{1},r_{1}}$) $\mathrm{V}^{\prime}(\mathrm{C_{i}})$ made from good cubes intersects the horizontal crossing of good cubes (of the strip $\mathrm{H}_{r_{1}}$) $\mathrm{H}^{\ell}_{r_{1}}$ in at least one good cube (this is true as the slice is 2-dimensional). Suppose without loss of generality that $\mathrm{V}_{1}^{\prime}(\mathrm{C_{i}})$ is intersecting this horizontal crossing. Concatenate then $\mathrm{V}_{1}^{\prime}(\mathrm{C_{i}})$ with $\mathrm{H}^{\ell}_{r_{1}}$ at the last good cube of $\mathrm{H}^{\ell}_{r_{1}}$ that intersects $\mathrm{V}_{1}^{\prime}(\mathrm{C_{i}})$; we denote the resulting path by $\mathcal{Q}_{1}$ and note that it is non-intersecting (see Figure 8, the bottom picture shows the final $\mathcal{Q}_{1}$). Now, observe that $(a_{3},b_{3})-(a_{2},b_{2})$ must be vertically oriented. We concatenate $\mathcal{Q}_{1}$ to the vertical crossing $\mathrm{V}_{a_{3}}^{\ell}$ (seen as path that goes from bottom to top) at the last cube of $\mathrm{V}_{a_{3}}^{\ell}$ that intersects $\mathcal{Q}_{1}$; we denote this path by $\mathcal{Q}_{2}$.
[…]
Figure 7. The path of smaller cubes is $\mathrm{V}(\mathrm{C_{i}})$. The path in yellow is $\mathrm{V}^{\prime}(\mathrm{C_{i}})$. In green and blue the two paths $\mathrm{V}^{\prime}_{1}(\mathrm{C_{i}})$ and $\mathrm{V}^{\prime}_{2}(\mathrm{C_{i}})$, respectively, obtained from $\mathrm{V}^{\prime}(\mathrm{C_{i}})$ that start from $\mathrm{C_{i}}$.
[…]
Figure 8. The gray path is $\mathrm{H}_{r_{1}}^{\ell}$. $\mathcal{Q}_{1}$ is the path that is left after concatenation in the picture below, adding the green and gray parts.
We continue in this way until we cross the midpoint (that is, we cross $y_{X}$). Here we continue interpreting the horizontal crossings as paths from left to right but change the orientation of the vertical crossings of good cubes to be from top to bottom. Furthermore, now instead of taking the $\ell$-th vertical crossing of good cubes we take the $(\lceil c\log n\rceil-\ell)$-th vertical crossing of good cubes to create the concatenations (the way of enumerating here is different so that that the concatenations are disjoint for different $\ell$ if we delete the good cubes from $\mathrm{V}^{\prime}(\mathrm{C_{i}})\cup\mathrm{V}^{\prime}(\mathrm{C_{j}})$, see Figure  9). To finish the construction, once we arrive at $\mathrm{T}_{s_{2},r_{2}}$, what is left is the concatenation to the vertical crossing of good cubes $\mathrm{V}^{\prime}(\mathrm{C_{j}})$. Here we proceed just as we did for $\mathrm{C_{i}}$. Select one path that starts from $\mathrm{C_{j}}$ and is made from good cubes of $\mathrm{V}^{\prime}(\mathrm{C_{j}})$ and intersects $\mathcal{Q}_{k-1}$ (the last concatenation available and that passes through $\mathrm{T}_{s_{2},r_{2}}$; there exists at least one such path of good cubes). Concatenate $\mathcal{Q}_{k-1}$ to the last path selected to obtain $\mathcal{Q}_{k}$. Note that the resulting path is self-avoiding by construction, and also note that for different $\ell$, the paths are disjoint if we delete the initial and final good cubes that come from $\mathrm{V}^{\prime}(\mathrm{C_{i}})$ and $\mathrm{V}^{\prime}(\mathrm{C_{j}})$. We denote $\mathcal{Q}_{k}$ by $\mathcal{Q}_{X}^{\ell}$. Delete from $\mathcal{Q}_{k}$ the good cubes that belong to $\mathrm{V}^{\prime}(\mathrm{C_{i}})$ and $\mathrm{V}^{\prime}(\mathrm{C_{j}})$, except for the cubes where the concatenations with $\mathrm{V}^{\prime}(\mathrm{C_{i}})$ and $\mathrm{V}^{\prime}(\mathrm{C_{j}})$ were made at the beginning and the end, and denote the resulting random path by $\mathcal{L}^{\ell}_{X}$ (see Figure  9). In particular, note that the path $\mathcal{L}_{X}^{\ell}$ is a sub-path of $\mathcal{Q}_{X}^{\ell}$ since the latter has $\mathrm{C_{i}}$ and $\mathrm{C_{j}}$ as endpoints (in Figure  9, $\mathcal{Q}_{X}^{\ell}$ corresponds to the blue path together with the green paths in Figure  9 while the path $\mathcal{L}_{X}^{\ell}$ is only the blue path). We point out that the construction of $\mathcal{Q}_{X}^{\ell}$ will be used only in the proof of the upper bound of the cover time above the connectivity threshold in Section 6 and Section  7.
[…]
Figure 9. The red path is $\mathcal{L}_{X}^{\ell_{1}}$, and the blue path is $\mathcal{L}_{X}^{\ell_{2}}$, they are disjoint as a consequence of the enumeration introduced for the long crossings. The green part (subset of $\mathrm{V}^{\prime}(\mathrm{C_{i}})\cup\mathrm{V}^{\prime}(\mathrm{C_{j}})$) allows us to reach $\mathrm{C_{i}}$ and $\mathrm{C_{j}}$.
5.2. Proof of the upper bound on the cover time with fixed radius
Theorem 5.1.
Let $d\geq 2$, let $G\in\mathcal{G}(n,r,d)$, and let $\varepsilon>0$ be arbitrarily small. Assume that $r\geq(1+\varepsilon)r_{g}$, and also assume $r=O(1)$. A.a.s., there is a constant $\gamma(d,r)>0$ such that
$\tau_{cov}(L_{1})\leq\gamma n\log^{2}n.$ (5.3)
Proof.
Since we use the construction given in Subsection 5.1, we recall our initial assumptions made on the good cubes $\mathrm{C_{i}}$ and $\mathrm{C_{j}}$: both cubes belong to the same strip and they both belong to parallelepipeds not at the boundary of the slice and separated by an even number of parallelepipeds. We will lift these assumptions later.
Consider any two vertices $x,y$ that belong to the good cubes $\mathrm{C_{i}}$ and $\mathrm{C_{j}}$ respectively, and both belong to the largest connected component of the renormalized graph of good cubes. As the good cubes $\mathrm{C_{i}}$ and $\mathrm{C_{j}}$ belong to long crossings of good cubes (in the renormalized graph) which at the same time contain long crossings of $G$ of length $\Theta(n^{1/d})$, by Theorem 2.13 $x,y$ then also belong to $L_{1}(G)$ a.a.s.
Select an induced path of $G\in\mathcal{G}(n,r,d)$ that starts at $x$, is contained in $\mathrm{V}_{1}^{\prime}(\mathrm{C_{i}})$ and has its final vertex at the last cube of $\mathrm{V}_{1}^{\prime}(\mathrm{C_{i}})$. Denote this path by $\mathcal{V}^{1}_{x}$ (note that this selection does not depend on the random variable $X$). We do the same with $\mathrm{V}_{2}^{\prime}(\mathrm{C_{i}})$ and denote the path by $\mathcal{V}^{2}_{x}$. We proceed in the same way for the vertex $y$.
Without loss of generality suppose that $\mathrm{V}_{1}^{\prime}(\mathrm{C_{i}})$ intersects $\mathcal{L}_{X}^{\ell}$. Denote by $\mathrm{C}^{1}$ the only cube of $\mathcal{L}_{X}^{\ell}$ that intersects $\mathrm{V}_{1}^{\prime}(\mathrm{C_{i}})$ (note that conditioned on $X$ the selection of $\mathrm{C}^{1}$ depends on $\ell$). Also suppose that $\mathrm{V}_{1}^{\prime}(\mathrm{C_{j}})$ intersects $\mathcal{L}_{X}^{\ell}$ in some good cube, and denote by $\mathrm{C}^{2}$ the first good cube of $\mathcal{L}_{X}^{\ell}$ that intersects $\mathrm{V}_{1}^{\prime}(\mathrm{C_{j}})$.
As $\mathcal{L}_{X}^{\ell}$ is a path of good cubes that begins at $\mathrm{C}^{1}$ and ends at $\mathrm{C}^{2}$, we may obtain an induced path (totally contained within $\mathcal{L}_{X}^{\ell}$) of $G\in\mathcal{G}(n,r,d)$ that starts at some vertex of $\mathcal{V}^{1}_{x}$ inside $\mathrm{C}^{1}$ and finishes at some vertex of $\mathcal{V}^{1}_{y}$ inside $\mathrm{C}^{2}$ (all cubes are good, and so we can use the giant components of the good cubes), and that intersects $\mathcal{V}^{1}_{x}$ and $\mathcal{V}^{1}_{y}$ only at its initial and final point, respectively. We denote this path by $J^{\ell}_{X}$, see Figure  10. We concatenate $\mathcal{V}^{1}_{x}$, $J^{\ell}_{X}$ and $\mathcal{V}^{1}_{y}$ in the obvious way (as Figure  10 suggests for the first part of the concatenation of blue and red). From this path, we may then extract an induced path with the same endpoints.
[…]
Figure 10. $\mathcal{V}^{1}_{x}$ and $J^{\ell}_{X}$ have exactly one vertex in common. We use this vertex to concatenate both paths.
We collect some properties of $\mathcal{J}_{X}^{\ell}$:
  1. (a)
    Let $\ell_{1}\neq\ell_{2}$. The set of vertices of $\mathcal{J}^{\ell_{1}}_{X}$ that are within good cubes that belong to $\mathcal{L}_{X}^{\ell_{1}}$ is disjoint from the set of vertices of $\mathcal{J}_{X}^{\ell_{2}}$ that are within the good cubes that belong to $\mathcal{L}_{X}^{\ell_{2}}$. If $\mathcal{J}_{X}^{\ell_{1}}$ and $\mathcal{J}_{X}^{\ell_{2}}$ have common vertices, these are inside of $\mathrm{V}^{\prime}(\mathrm{C_{i}})$ or $\mathrm{V}^{\prime}(\mathrm{C_{j}})$.
  2. (b)
    The number of edges that have some endpoint inside of $\mathrm{V}^{\prime}(\mathrm{C_{i}})\cup\mathrm{V}^{\prime}(\mathrm{C_{j}})$ and belong to some path $\mathcal{J}_{X}^{\ell}$ is $\Theta(\log n)$.
For point (a) recall that if $\ell_{1}\neq\ell_{2}$, then $\mathcal{L}_{X}^{\ell_{1}}$ does not intersect $\mathcal{L}_{X}^{\ell_{2}}$, and as $\mathcal{J}^{\ell}_{X}$ is totally contained within $\mathcal{L}_{X}^{\ell}\cup\mathrm{V}^{\prime}(\mathrm{C_{i}})\cup\mathrm{V}^{\prime}(\mathrm{C_{j}})$, part (a) follows.
For (b), recall that $\mathrm{V}^{\prime}(\mathrm{C_{i}})$ has logarithmic length (in the renormalized lattice where we see cubes as points of $\mathbb{Z}^{d}$). Since the path $\mathcal{V}^{1}_{x}$ is chosen to be an induced path it has a logarithmic number of edges. Note that the other edges of $\mathcal{J}_{X}^{\ell}$ with at least one endpoint that is inside of $\mathrm{V}^{\prime}(\mathrm{C_{i}})$ must have this endpoint inside of $\mathrm{C}^{1}$. Since the path is induced and the cube $\mathrm{C}^{1}$ has constant size, the number of edges from some realization $\mathcal{J}_{X}^{\ell}$ and with at least one endpoint that is inside of $\mathrm{C}^{1}$ is $O(1)$. Finally, since there is a logarithmic number of such cubes $\mathrm{C}^{1}$ (one for each $\ell$) the number of such vertices is logarithmic. The same is true for $\mathrm{V}^{\prime}(\mathrm{C_{j}})$ and we obtain property (b).
Now, let $Y$ be a discrete uniform random variable with values in $\{1,...,\lfloor c\log n\rfloor\}$ and independent from $X$. Define the random path $\mathcal{J}_{X}^{Y}$ and the function
$\psi(u,v)=\begin{cases}\text{1},&\text{if $\{u,v\}\in\mathcal{J}_{X}^{Y}$ and the path visits $u$ first, }\\ \text{-1},&\text{if $\{u,v\}\in\mathcal{J}_{X}^{Y}$ and the path visits $v$ first,}\\ \text{0},&\text{otherwise.}\end{cases}$
We define a unit flow $\theta$ between $x$ and $y$ by
$\theta(u,v)=\mathbb{E}(\psi(u,v))$ (5.4)
for $u,v\in G$, where the expectation is with respect to $X,Y$. Note that
$|\theta(u,v)|\leq\mathbb{P}_{X,Y}(\{u,v\}\in\mathcal{J}_{X}^{Y})=:p(u,v)$ (5.5)
We next group the parallelepipeds $\mathrm{T}$ that contain some realization of $\mathcal{J}^{Y}_{X}$ according to their level $t$ as follows (see Figure 11):
$D_{t}=\begin{cases}\{\mathrm{T}_{i,j}:i=s_{1}+t,\,r_{1}\leq j\leq r_{1}+t\},&\text{if $0\leq t\leq\frac{s_{2}-s_{1}+1}{2}$, }\\ \{\mathrm{T}_{i,j}:i=s_{1}+t,\,r_{1}\leq j\leq r_{1}+s_{2}-s_{1}-t-1\},&\text{if $\frac{s_{2}-s_{1}+1}{2}+1$}\leq t\leq s_{2}-s_{1}+1.\end{cases}$
[…]
Figure 11. Parallelepipeds at level $t$ (i.e., their projections onto the first $2$ dimensions). See the projected probability over the right line of $D_{t}$ inside the blue dotted lines.
$x$$y$$z$
Figure 12. Using triangle inequality for vertices in different strips and in the same 2-dimensional slice.
We are now ready to bound the effective resistance between $x,y$. Recall that $\mathfrak{r}(e)=1$, and so, since $\theta^{2}(u,v)\leq p^{2}(u,v)$, by Definition 2.1, in order to bound $\mathcal{R}(x\longleftrightarrow y)$, it suffices to bound $\sum_{e\in E}p^{2}(e)$. More precisely, we may split the edge set according to levels of parallelepipeds and get:
$\mathcal{R}(x\longleftrightarrow y)\leq\sum_{u\in\mathrm{V}^{\prime}(\mathrm{C_{i}})\cup\mathrm{V}^{\prime}(\mathrm{C_{j}})}\sum_{\begin{subarray}{c}v\in L_{1}\\ v\sim u\end{subarray}}p^{2}(u,v)+\sum_{t}\sum_{\mathrm{T}_{i,j}\in D_{t}}\sum_{\begin{subarray}{c}u\in\mathrm{T}_{i,j}\\ u\notin\mathrm{V}^{\prime}(\mathrm{C_{i}})\cup\mathrm{V}^{\prime}(\mathrm{C_{j}})\end{subarray}}\sum_{\begin{subarray}{c}v\in L_{1}\\ v\sim u\end{subarray}}p^{2}(u,v).$ (5.6)
The first double sum on the right hand side can be bounded by $O(\log n)$ as the number of edges that have positive probability of belonging to the path is $\Theta(\log n)$ (see properties of $\mathcal{J}^{\ell}_{X}$ above).
Consider then the remaining term. Let $u\in\mathrm{T}_{i,j}$ with $\mathrm{T}_{i,j}\in D_{t}$ and $\{u,v\}\in E$ and $u\notin\mathrm{V}^{\prime}(\mathrm{C_{i}})\cup\mathrm{V}^{\prime}(\mathrm{C_{j}})$. We now claim that the probability that $\mathrm{T}_{i,j}\in D_{t}$, with $1\leq t\leq\frac{s_{2}-s_{1}+1}{2}$, is part of the path of parallelepipeds $\mathcal{L}_{X}$ is bounded above by $\frac{c^{\prime}}{t}$ for some constant $c^{\prime}$ that can be chosen independent of $n$ and $t$ (for $t=0$ we bound this probability by 1). To show this, denote by $y_{X}^{t}$ the intersection of the segment between the left bottom corner point of $\mathrm{T}_{s_{1},r_{1}}$ and $y_{X}$ and the left line of level $t$ (see Figure  11). By similarity of triangles, $y_{X}^{t}$ is uniform over the left line of level $t$ (see for example the triangles $Ay^{t}_{X}B$ and $Ay_{X}C$ in Figure 11). Now, observe that in order for $\mathrm{T}_{i,j}\in D_{t}$ to be part of $\mathcal{L}_{X}$ the random point $y_{X}^{t}$ has to belong to the left face of $\mathrm{T}_{i,j}$ or to $\mathrm{T}_{i,j-1}$ (the left face of the parallelepiped immediately below, in case such parallelepiped exists): indeed, if $y_{X}^{t}$ were above $\mathrm{T}_{i,j}$, then the parallelepiped does not belong to the path, and if $y^{t}_{X}$ were below $\mathrm{T}_{i,j-1}$, as the angle of the red line is at most $\pi/4$, the red line in Figure 11 does not intersect $\mathrm{T}_{i,j}$. Therefore, the probability that $\mathrm{T}_{i,j}\in D_{t}$ belongs to $\mathcal{L}_{X}$ is bounded from above by the probability that $y_{X}^{t}$ belongs to the left faces of two cubes at level $t$. Since $y_{X}^{t}$ is uniform in its range, then this probability is at most $\frac{2}{t}$. By symmetry, the same is true for $\frac{s_{1}-s_{1}+1}{2}+1\leq t\leq s_{2}-s_{1}$ (for $t=s_{2}-s_{1}+1$ we bound this probability by $1$); in this case we obtain an upper bound of $\frac{2}{s_{2}-s_{1}-t-1}$ for the probability that $\mathrm{T}_{i,j}\in D_{t}$.
Consider a fixed realization of $\mathcal{L}_{X}$, and suppose first that $\{u,v\}\in E$ such that $u$ belongs to some parallelepiped $\mathrm{T}_{i,j}$ of $\mathcal{L}_{X}$. We may also assume that the vertex $u$ does not belong to the good cubes of $\mathrm{V}^{\prime}(\mathrm{C_{i}})\cup\mathrm{V}^{\prime}(\mathrm{C_{j}})$ (otherwise the contribution is already counted in the first double sum). Then, recall that $\{u,v\}$ belongs to at most one path of the form $\mathcal{J}_{X}^{Y}$ (by property (a) of the path $\mathcal{J}_{X}^{\ell}$ above). By independence of $X,Y$ and as $\{u,v\}$ belongs to one unique path $\mathcal{J}_{X}^{Y}$ (with $X$ fixed), and as the path is chosen uniformly from a set of $\lfloor c\log n\rfloor$ possible realizations, we have
$\mathbb{P}_{X,Y}(\{u,v\}\in\mathcal{J}^{Y}_{X}|X)\leq\frac{1}{\lfloor c\log n\rfloor}.$ (5.7)
Furthermore, recall that $u$ belongs to the parallelepiped of $\mathrm{T}_{i,j}\in D_{t}$. Thus, by the above, since the choice of the parallelepiped in $D_{t}$ (with $1\leq t\leq\frac{s_{2}-s_{1}+1}{2}$) and the choice of the crossing are independent, $p(u,v)\leq\frac{c^{\prime\prime}}{t\log n}$ for some $c^{\prime\prime}>0$. The same occurs by symmetry for the other levels.
Fix then a parallelepiped $\mathrm{T}$ that belongs to some level $t$. Recall that the volume of the parallelepiped $\mathrm{T}$ is $\Theta(\log^{2}n)$, as in 2 dimensions its side length is $Mr\lceil\alpha\log n\rceil$ and it is $Mr$ in the other dimensions (recall that $r$ is assumed to be constant in this section). Hence, the number of edges that have at least one endpoint inside of $\mathrm{T}$ is by Corollary  2.9 a.a.s. $O(\log^{2}n)$. We substitute these observations into the right hand expression of (5.6) and obtain
$\sum_{1\leq t\leq\frac{s_{2}-s_{1}+1}{2}}\sum_{\mathrm{T}_{i,j}\in D_{t}}\sum_{\begin{subarray}{c}u\in\mathrm{T}_{i,j}\\ u\notin\mathrm{V}^{\prime}(\mathrm{C_{i}})\cup\mathrm{V}^{\prime}(\mathrm{C_{j}})\end{subarray}}\sum_{v\in L_{1}}p^{2}(u,v)$ $\leq\sum_{1\leq t\leq\frac{s_{2}-s_{1}+1}{2}}\sum_{\mathrm{T}_{i,j}\in D_{t}}\sum_{\begin{subarray}{c}u\in\mathrm{T}_{i,j}\\ u\notin\mathrm{V}^{\prime}(\mathrm{C_{i}})\cup\mathrm{V}^{\prime}(\mathrm{C_{j}})\end{subarray}}\sum_{v\in L_{1}}\frac{c^{\prime\prime\prime}}{t^{2}\log^{2}n}$ (5.8)
$\leq\sum_{1\leq t\leq\frac{s_{2}-s_{1}+1}{2}}\sum_{\mathrm{T}_{i,j}\in D_{t}}\log^{2}n\frac{c^{\prime\prime\prime\prime}}{t^{2}\log^{2}n}$
$\leq\sum_{1\leq t\leq\frac{s_{2}-s_{1}+1}{2}}(t+1)\log^{2}n\frac{c^{\prime\prime\prime\prime}}{t^{2}\log^{2}n}$
$=\sum_{1\leq t\leq\frac{s_{2}-s_{1}+1}{2}}O\left(\frac{1}{t}\right)=O(\log n).$
It is clear that the contribution of $D_{0}$ to the inequality is $O(1)$. By symmetry around $y_{X}$, one obtains the same bound for the remaining levels $\frac{s_{2}-s_{1}+1}{2}<t\leq s_{2}-s_{1}+1$. Then $\mathcal{R}(x\longleftrightarrow y)=O(\log n)$.
We now lift the assumptions on the parallelepipeds: consider first two vertices $x,y$ of $L_{1}$ that are in the same strip of a 2-dimensional slice, but now separated by an odd number of parallelepipeds that are not at the boundary. Let $x\in\mathrm{T}_{s_{1}^{\prime},r_{1}}$ and $y\in\mathrm{T}_{s_{2}^{\prime},r_{1}}$ with $s_{2}^{\prime}-s_{1}^{\prime}$ even such that neither parallelepiped is at the boundary. By the construction given above there exists $x^{\prime}\in L_{1}(G)$ with $x^{\prime}\in\mathrm{T}_{s_{1}^{\prime}+1,r_{1}}$ such that $\mathcal{R}(x^{\prime}\longleftrightarrow y)=O(\log n)$. Note that $d_{E}(x,x^{\prime})=O(\log n)$, since $x$ and $x^{\prime}$ are in adjacent parallelepipeds, and any parallelepiped has logarithmic side length in the first 2 dimensions and fixed side length in the other dimensions. By Corollary  2.12, a.a.s., the graph distance between $x$ and $x^{\prime}$ is then also $O(\log n)$, and thus a.a.s. $\mathcal{R}(x\longleftrightarrow x^{\prime})=O(\log n)$. Then, by application of the triangle inequality (see Lemma 2.4), also a.a.s. $\mathcal{R}(x\longleftrightarrow y)=O(\log n)$. Also, since any two vertices in the same parallelepiped (not necessarily in good cubes, but still in $L_{1}(G)$) satisfy that their graph distance is $O(\log n)$ (see Corollary 2.12, the largest side length of a parallelepiped is $O(\log n)$), by another application of the triangle inequality, also a.a.s. the effective resistance between them is $O(\log n)$. Hence, for any $x^{\prime},y^{\prime}\in L_{1}(G)$ belonging to the largest connected component of the renormalized graph inside the same strip of a $2$-dimensional slice not at the boundary, we have
$\mathcal{R}(x\longleftrightarrow y)=O(\log n).$ (5.9)
By the same idea, we may extend our bounds to vertices in parallelepipeds of a strip on the boundary of the 2-dimensional slice, as for these vertices there exists a vertex at graph distance $O(\log n)$ in a parallelepiped not at the boundary.
We now extend the result to any two vertices $x,y$ of $L_{1}(G)$ in the same 2-dimensional slice: to do so, suppose that $x,y\in\mathcal{K}^{2}_{n}$. Let $x\in\mathrm{H}_{r}$ ($x$ is in the $r$-the horizontal strip) and let $y\in\mathrm{V}_{s}$ ($y$ in the $s$-th vertical strip). Note that this election is always possible as any vertex belongs to one vertical strip and one horizontal strip. Pick $z\in L_{1}$ such that $z\in\mathrm{V}_{s}\cap\mathrm{H}_{r}=\mathrm{T}_{s,r}$ (see Figure  12). Note that such $z$ must exist as $\mathrm{T}_{s,r}$ is the intersection of two strips that contain long crossings of good cubes, which at the same time contain long crossings of $L_{1}$. As $x$ and $y$ are in the same strip, by the above, a.a.s. $\mathcal{R}(x\longleftrightarrow z)=O(\log n)$. Similarly, a.a.s. $\mathcal{R}(y\longleftrightarrow z)=O(\log n)$, and applying again the triangle inequality the result holds as well.
Now, in general, suppose that there is no 2-dimensional slice containing $x,y$ (while always a plane containing $x,y$ exists, we cannot ensure the existence of a slice, which by our definition is assumed to be axis-parallel). Note that in case two different slices have non-empty intersection, this intersection is a line of cubes such that only one coordinate is not fixed. In this case, as before, there exists at least a vertex $z\in G\in\mathcal{G}(n,r,d)$ that belongs to $L_{1}$ and that is in the intersection of both slices. Then, for any two vertices $x^{\prime},y^{\prime}$ such that $x^{\prime}$ is in the first slice and $y^{\prime}$ is in the second slice, by the above, a.a.s. $\mathcal{R}(x^{\prime}\longleftrightarrow z)=O(\log n)$ and $\mathcal{R}(y^{\prime}\longleftrightarrow z)=O(\log n)$, and thus, by the triangle inequality, also $\mathcal{R}(x^{\prime}\longleftrightarrow y^{\prime})=O(\log n)$. To conclude, observe that for any two vertices $x,y\in G$, there exists a sequence of slices $(\mathcal{K}_{n}^{2})_{1},...,(\mathcal{K}_{n}^{2})_{s}$, with $s\leq d-1$, such that for any $11\leq i\leq s-$ 1, $(\mathcal{K}_{n}^{2})_{s}$ has non-empty intersection with $(\mathcal{K}_{n}^{2})_{s+1}$, and moreover, $x\in(\mathcal{K}_{n}^{2})_{1}$, and $y\in(\mathcal{K}_{n}^{2})_{s}$. By applying the triangle inequality iteratively, we then have a.a.s. $\mathcal{R}(x\longleftrightarrow y)=O(\log n)$.
By this result, together with Lemma 2.10 we then immediately conclude that a.a.s.
$\tau_{cov}\leq\gamma n\log^{2}n,$ (5.10)
for some $\gamma(d,r)>0$ (uniformly bounded from below), and the theorem follows. ∎
6. Upper bound for non-fixed radius
In this section we prove the upperbound for RGGs with radii tending to infinity but below the threshold of connectivity. We begin by noting that the proof given in Section 5 is not enough, as it gives as an upper bound on the resistance of $O(\log n)$, whereas we need an upper bound of magnitude $O(\log(n)/r^{d})$. With this in mind, we therefore introduce a new construction inspired by the one in Section 5 but with refinements that allow for a better flow dissipation. This structure will then also be used in the proof of the upper bound just above the connectivity threshold.
6.1. Resistance bound in the backbone of dense cubes
Recall the definition of $\mu$ introduced in the tessellation of cubes $(\mathrm{Q_{i}})$ in Section 4 of cubes of side length $r/\mu$. In this section, we choose $\mu(n)=\mu+O(1/n)$ to be the smallest real number larger or equal to $\mu$ such that any cube $\mathrm{Q}$ is either completely contained in $\Lambda_{n}$ or it shares at most some faces, so that we avoid boundary issues. With abuse of notation, we denote below $\mu(n)=\mu$. Let $\beta$ be a sufficiently small constant. We say that $\mathrm{Q_{i}}$ is dense if it has at least $\beta r^{d}$ vertices of $G\in\mathcal{G}(n,r,d)$, otherwise we say that it is non-dense.
We see cubes $\mathrm{Q}$ as sites of $\mathbb{Z}^{d}$. By Chernoff’s inequality, the probability that a cube is non-dense is bounded from above by $e^{-\kappa r^{d}}$ for some $\kappa=\kappa(\beta)>0$ (and $\beta$ small enough). We note that $\kappa$ is a fixed constant and does not depend on $r$. Considering dense cubes as occupied sites of a Bernoulli site percolation model, for $r\geq r_{0}$ with $r_{0}$ large enough, we are thus in the super-critical phase of Bernoulli site percolation. That is, a.a.s., there exists a giant component corresponding to dense cubes. We denote by $\mathcal{C}^{site}_{\mathrm{Q}}(n)$ the giant (biggest) cluster of the associated Bernoulli process of dense cubes $\mathrm{Q}$. We prove the following regarding the maximum of the effective resistance in the giant cluster:
Theorem 6.1.
Let $r\geq r_{0}$ for a sufficiently large constant $r_{0}>0$ and let also $r=O(r_{c})$. Let $x,y\in G\in\mathcal{G}(n,r,d)$ such that $x\in\mathrm{Q_{i}}$ and $y\in\mathrm{Q_{j}}$ and $\mathrm{Q_{i}},\mathrm{Q_{j}}\in\mathcal{C}^{site}_{\mathrm{Q}}(n)$. Then, a.a.s.,
$\mathcal{R}(x\longleftrightarrow y)=O\left(\frac{\log n}{r^{2d}}\right).$
Proof.
First, make $C$ big enough such that $\beta r^{d}$ is at least 3 and that the dense cubes dominate a supercritical Bernoulli site percolation process. Divide $\Lambda_{n}$ into disjoint $\alpha$-logarithmic strips of cubes $\mathrm{Q_{i}}$. We use the structure introduced in Subsection 5.1, with the difference that we consider cubes $\mathrm{Q_{i}}$ instead of cubes $\mathrm{C_{i}}$, and dense cubes play the role of good cubes. Recall that we obtain the construction given there a.a.s.
In the following we assume vertices $x,y$ belonging to a fixed $2$-dimensional slice (we generalize this later). Let $\mathcal{Q}^{Y}_{X}$ (here $Y$ is a uniform discrete random variable independent from $X$ with values in $\{1,\dots,\lfloor c\log n\rfloor\}$, see the construction of Subsection 5.1 for the definition of $c$ and Theorem 5.1 for details of the construction of $\mathcal{Q}^{\ell}_{X}$ with a fixed value $\ell$ of $Y$) be the random path of dense cubes between two fixed dense cubes $\mathrm{Q_{i}}$ and $\mathrm{Q_{j}}$, such that both belong to the giant component of dense cubes in $\mathcal{C}^{site}_{\mathrm{Q}}(n)$, and as in Subsection 5.1, we may assume that the two dense cubes belong to some parallelepipeds $\mathrm{Q_{i}}\in\mathrm{T}_{s_{1},r_{1}}$ and $\mathrm{Q_{j}}\in\mathrm{T}_{s_{2},r_{2}}$ such that $r_{1}=r_{2}$, $s_{2}>s_{1}$, $s_{2}-s_{1}$ is odd, and that both parallelepipeds do not belong to a strip that is at the boundary of the 2-dimensional slice. Pick vertices $x,y$ of $G\in\mathcal{G}(n,r,d)$ such that $x\in\mathrm{Q_{i}}$ and $y\in\mathrm{Q_{j}}$. For each dense cube that is part of some realization of $\mathcal{Q}^{Y}_{X}$ select $\lfloor\beta r^{d}\rfloor$ vertices, except for $\mathrm{Q_{i}}$ and $\mathrm{Q_{j}}$ where we may choose $\lfloor\beta r^{2}\rfloor-1$ vertices distinct from $x,y$ (or $\lfloor\beta r^{2}\rfloor-2$ in case $\mathrm{Q_{i}}=\mathrm{Q_{j}}$).
Conditional on $X$ and $Y$, we construct a random path of $G\in\mathcal{G}(n,r,d)$ that starts at $x$, finishes at $y$, and is entirely contained within the path of dense cubes $\mathcal{Q}_{X}^{Y}$ as follows. First, select uniformly at random and independently from $X,Y$ one of the $\lfloor\beta r^{d}\rfloor-1$ vertices that belong to the dense cube $\mathrm{Q_{i}}$ (one out of the $\lfloor\beta r^{d}\rfloor-1$ vertices previously selected), and connect it by an edge to the vertex $x$: this is the first edge of our path. For subsequent steps we do something similar: suppose that the last vertex that we chose belongs to the dense cube $\mathrm{Q}_{s}$. We then choose uniformly and independently at random from the $\lfloor\beta r^{d}\rfloor$ vertices that belong to the dense cube that follows in the path $\mathcal{Q}^{Y}_{X}$. We proceed until the last selected vertex belongs to the cube just before $\mathrm{Q_{j}}$ (the latter is thee dense cube that contains $y$). In this case we select uniformly at random between the $\lfloor\beta r^{d}\rfloor-1$ selected vertices in $\mathrm{Q_{j}}$ and finally connect all these last vertices by an edge to $y$ (if we had $\mathrm{Q_{i}}=\mathrm{Q_{j}}$, then we would pick $\lfloor\beta r^{d}\rfloor-2$ vertices inside the cube and form paths of length $2$ from $x$ to $y$ through these vertices). This process gives rise to $\lfloor\beta r^{d}\rfloor-2$ internally vertex-disjoint random paths of $G\in\mathcal{G}(n,r,d)$ with initial vertex $x$ and final vertex $y$. We denote this random path of $G\in\mathcal{G}(n,r,d)$ by $\mathcal{J}_{X,Z}^{Y}$ and denote its law by $\mathbb{P}_{X,Y,Z}$, where $Z$ denotes the randomness that comes from the path selection conditioned on $X,Y$. Observe that the number of edges that have some endpoint inside of $\mathrm{V}^{\prime}(\mathrm{Q_{i}})\cup\mathrm{V}^{\prime}(\mathrm{Q_{j}})$ is $O(r^{2d}\log n)$, since $\mathrm{V}^{\prime}(\mathrm{Q_{i}})$ (see also Subsection  5.1) is made of $\Theta(\log n)$ dense cubes, each of which has $O(r^{d})$ selected vertices, and any such vertex has $\Theta(r^{d})$ neighbors among selected vertices inside its own cube and inside a constant number of cubes around it. Since between any two consecutive dense cubes there are $\Theta(r^{2d})$ edges, and we choose uniformly at random one of them, we obtain in the case $u\neq x,u\neq y,v\neq x,v\neq y$,
$\mathbb{P}_{X,Y,Z}(\{u,v\}\in\mathcal{J}_{Z}^{Y}|X,Y)=O\left(\frac{1}{r^{2d}}\right),$
whereas in the case |$|\{u,v\}\cap\{x,y\}|=1$, we have
$\mathbb{P}_{X,Y,Z}(\{u,v\}\in\mathcal{J}_{X,Z}^{Y}|X,Y)=O\left(\frac{1}{r^{d}}\right).$
As before, denote by $p(u,v)$ the probability that the edge $\{u,v\}$ belongs to the random path $\mathcal{J}_{X,Z}^{Y}$. Using the random path $\mathcal{J}_{X,Y}^{Y}$, we define a flow in the same way as in the proof of Theorem  5.1, and we obtain
$\mathcal{R}(x\longleftrightarrow y)\leq$ $\sum_{\begin{subarray}{c}v\in L_{1}\\ y\sim v\end{subarray}}p^{2}(y,v)+\sum_{\begin{subarray}{c}v\in L_{1}\\ x\sim v\end{subarray}}p^{2}(x,v)$
$+\sum_{\begin{subarray}{c}u\in\mathrm{V}^{\prime}(\mathrm{Q}_{i})\cup\mathrm{V}^{\prime}(\mathrm{Q}_{j})\\ u\neq x,u\neq y\end{subarray}}\sum_{\begin{subarray}{c}v\in L_{1}\\ v\neq x,v\neq y\\ u\sim v\end{subarray}}p^{2}(u,v)$
$+\sum_{t}\sum_{\mathrm{T}_{i,j}\in D_{t}}\sum_{\begin{subarray}{c}u\in\mathrm{T}_{i,j}\\ u\notin\mathrm{V}^{\prime}(\mathrm{Q}_{i})\cup\mathrm{V}^{\prime}(\mathrm{Q}_{j})\\ u\neq x,u\neq y\end{subarray}}\sum_{\begin{subarray}{c}v\in L_{1}\\ u\sim v\\ v\neq x,v\neq y\end{subarray}}p^{2}(u,v).$ (6.1)
To bound the first two sums, since $p(x,v)=O\left(\frac{1}{r^{d}}\right)$, and since $x$ and $y$ send flow to different $O(r^{d})$ vertices, we have $\sum_{v\in L_{1}}\big{(}p^{2}(y,v)+p^{2}(x,v)\big{)}=O\left(1/r^{d}\right)$. For the sum in the second row, since the number of edges with at least one endpoint inside of $\mathrm{V}^{\prime}(\mathrm{Q_{i}})\cup\mathrm{V}^{\prime}(\mathrm{Q_{j}})$ is $O(r^{2d}\log n)$, and since $p(u,v)=O(1/r^{2d})$,
$\sum_{\begin{subarray}{c}u\in\mathrm{V}^{\prime}(\mathrm{Q}_{i})\cup\mathrm{V}^{\prime}(\mathrm{Q}_{j})\\ u\neq x,u\neq y\\ \end{subarray}}\sum_{\begin{subarray}{c}v\in L_{1}\\ v\neq x,v\neq y\\ u\sim v\end{subarray}}p^{2}(u,v)=O\left(\frac{\log n}{r^{2d}}\right).$
To bound the term in the last row, recall that, as in the proof of Theorem 5.1 in Section 5, the probability that a dense cube inside of some parallelepiped $\mathrm{T}\in D_{t}$ belongs to $\mathcal{Q}_{X}^{Y}$ is $O\left(\frac{1}{t\log n}\right)$ for $1\leq t\leq\frac{s_{2}-s_{1}+1}{2}$ and $O\left(\frac{1}{\log n}\right)$ for $t=0$ (if the parallelepiped does not belong to $\mathrm{V}^{\prime}(\mathrm{Q_{i}})\cup\mathrm{V}^{\prime}(\mathrm{Q_{j}})$), since there are $\Omega(t)$ parallelepipeds at level $t\geq 1$, and we choose one of the $\Omega(\log n)$ crossings of dense cubes, uniformly at random. Now, observe that we have the additional randomness coming from the fact for any pair of consecutive dense cubes, we choose u.a.r. a vertex inside each cube, giving another factor $O(1/r^{2d})$, as mentioned above. All three choices are independent of each other, and thus, for edges $\{u,v\}$ appearing in the last term, we have
$p(u,v)=O\left(\frac{1}{tr^{2d}\log n}\right).$
The number of edges with positive probability $p$ and with at least one endpoint inside of a parallelepiped $\mathrm{T}$ is a.a.s. $O(r^{2d}\log^{2}n)$, as there are $O(\log^{2}n)$ dense cubes inside $\mathrm{T}$ (the volume of $\mathrm{T}$ is $\Theta(r^{d}\log^{2}n)$, and each cube has volume $\Theta(r^{d})$), each dense cube contains $O(r^{d})$ vertices, and the selected vertices are connected by an edge to another $O(r^{d})$ selected vertices by construction. Thus
$\sum_{1\leq t\leq\frac{s_{2}-s_{1}+1}{2}}$ $\sum_{\mathrm{T}_{i,j}\in D_{t}}\sum_{\begin{subarray}{c}u\in\mathrm{T}_{i,j}\\ u\notin\mathrm{V}^{\prime}(\mathrm{Q}_{i})\cup\mathrm{V}^{\prime}(\mathrm{Q}_{j})\\ u\neq x,u\neq y\end{subarray}}\sum_{\begin{subarray}{c}v\in L_{1}\\ u\sim v\\ v\neq x,v\neq y\end{subarray}}p^{2}(u,v)$ (6.2)
$\leq\sum_{1\leq t\leq\frac{s_{2}-s_{1}+1}{2}}\sum_{\mathrm{T}_{i,j}\in D_{t}}\sum_{\begin{subarray}{c}u\in\mathrm{T}_{i,j}\\ u\notin\mathrm{V}^{\prime}(\mathrm{Q}_{i})\cup\mathrm{V}^{\prime}(\mathrm{Q}_{j})\\ u\neq x,u\neq y\end{subarray}}\sum_{\begin{subarray}{c}v\in L_{1}\\ u\sim v\\ v\neq x,v\neq y\end{subarray}}\frac{c^{\prime\prime\prime}}{t^{2}r^{4d}\log^{2}n}$
$\leq\sum_{1\leq t\leq\frac{s_{2}-s_{1}+1}{2}}\sum_{\mathrm{T}_{i,j}\in D_{t}}r^{2d}\log^{2}n\frac{c^{\prime\prime\prime\prime}}{t^{2}r^{4d}\log^{2}n}$
$\leq\sum_{1\leq t\leq\frac{s_{2}-s_{1}+1}{2}}(t+1)\frac{c^{\prime\prime\prime\prime}}{t^{2}r^{2d}}$
$=O\left(\frac{1}{r^{2d}}\right)\sum_{1\leq t\leq\frac{s_{2}-s_{1}+1}{2}}O\left(\frac{1}{t}\right)$
$=O\left(\frac{\log n}{r^{2d}}\right).$
It is clear that the contribution of $D_{0}$ to the formula is $O\left(\frac{1}{r^{2d}}\right)$. By symmetry, the contribution of $D_{t}$ with $\frac{s_{2}-s_{1}}{2}<t\leq s_{2}-s_{1}+1$ is also $O\left(\frac{1}{tr^{2d}}\right)$, and thus
$\mathcal{R}(x\longleftrightarrow y)=O\left(\frac{\log n}{r^{2d}}\right).$
Recall that we assumed that $x,y$ belong to some dense cubes $\mathrm{Q_{i}}$ and $\mathrm{Q_{j}}$, that at the same time are inside of some parallelepipeds in the same horizontal strip and separated by an even number of parallelepipeds. Before we lift these conditions, we first provide a useful construction: we claim that if $x^{\prime},y^{\prime}\in L_{1}(G)$ with $x^{\prime}\in\mathrm{Q_{i}}^{\prime}$ and $y^{\prime}\in\mathrm{Q_{j}}^{\prime}$ with both cubes being dense, and such that $\mathrm{Q_{i}}^{\prime}$ and $\mathrm{Q_{j}}^{\prime}$ are connected by a path $\mathcal{L}$ of dense cubes of length $O(\log n)$, then $\mathcal{R}(x^{\prime}\longleftrightarrow y^{\prime})=\left(\frac{\log n}{r^{2d}}\right)$.
To see this, let $x^{\prime}$ send an amount of $\frac{1}{\lfloor\beta r^{d}\rfloor-1}$ flow to the $\lfloor\beta r^{d}\rfloor-1$ previously selected vertices of $\mathrm{Q_{i}}^{\prime}$. From there on, let the previously selected vertices to send a uniformly distributed flow to the selected $\lfloor\beta r^{d}\rfloor$ vertices that belong to the dense cube that follows in the path that connects $\mathrm{Q_{i}}^{\prime}$ and $\mathrm{Q_{j}}^{\prime}$. We continue distributing in a uniform way the flow until we arrive until $\mathrm{Q_{j}}^{\prime}$. Once the flow reached the $\lfloor\beta r^{d}\rfloor-1$ selected vertices of $\mathrm{Q_{j}}^{\prime}$, send all the flow back to $y^{\prime}$. Denote this flow by $\theta$. We then have
$\mathcal{R}(x^{\prime}\longleftrightarrow y^{\prime})\leq\sum_{\begin{subarray}{c}z\in\mathrm{Q_{i}}^{\prime}\\ x^{\prime}\sim z\end{subarray}}\theta^{2}(x^{\prime},z)+\sum_{\begin{subarray}{c}z\in\mathrm{Q_{j}}^{\prime}\\ y^{\prime}\sim z\end{subarray}}\theta^{2}(y^{\prime},z)+\sum_{\mathrm{Q}\in\mathcal{L}}\sum_{\begin{subarray}{c}u\in\mathrm{Q}\\ v\in L_{1}\\ u\sim v\end{subarray}}\theta^{2}(u,v)$ (6.3)
As before, the first two sums contribute $O(r^{d})O(1/r^{2d})=O(1/r^{d})$. For the third sum, note that there are $O(\log n)$ dense cubes in $\mathcal{L}$, and between two consecutive dense cubes there are $O(r^{2d})$ edges with flow $O\left(\frac{1}{r^{2d}}\right)$ on each of them. Thus
$\mathcal{R}(x^{\prime}\longleftrightarrow y^{\prime})=O\left(\frac{1}{r^{d}}\right)+\sum_{|\mathcal{L}|}O\left(r^{2d}\right)O\left(\frac{1}{r^{4d}}\right)=O\left(\frac{\log n}{r^{2d}}\right),$
and the claim is proved.
We now show how this claim allows us to prove the bound on the effective resistance for any $x\in\mathrm{C_{i}}^{\prime}\in\mathrm{T}_{s_{1},r_{1}}$ and $y\in\mathrm{C_{j}}^{\prime}\in\mathrm{T}_{s_{2},r_{2}}$ in the same strip, first assuming that both parallelepipeds do not belong to strips on the boundary. First, if $x$ and $y$ are separated by an odd number of parallelepipeds (that is, $s_{2}-s_{1}$ even), we find $x^{\prime}$ in a parallelepiped adjacent to $x$, at Euclidean distance (in the renormalized lattice of cubes) $O(\log n)$ from $x$. Thus, by the claim, $\mathcal{R}(x\longleftrightarrow x^{\prime})=O\left(\frac{\log n}{r^{2d}}\right)$, and since $x^{\prime}$ and $y$ are separated by an even number of parallelepipeds, by the above, $\mathcal{R}(x^{\prime}\longleftrightarrow y)=O\left(\frac{\log n}{r^{2d}}\right)$, and finally by Lemma 2.4, also $\mathcal{R}(x\longleftrightarrow y)=O\left(\frac{\log n}{r^{2d}}\right)$. As before, by the same idea, we may extend our bounds also to arbitrary vertices inside the same parallelepiped (in cubes not satisfying all restrictions) and to vertices in parallelepipeds of a strip on the boundary.
To generalize the inequality between any pair of strips we use the same idea as in Section 5 from Figure 12: for any pair of strips, we find another vertex $z$ in the intersection of the strips that contain $x$ and $y$ and that intersect at some parallelepiped $\mathrm{T}$ that contains the vertex $z$ such that $\mathcal{R}(x\longleftrightarrow z)$ and $\mathcal{R}(y\longleftrightarrow z)$ are $O\left(\frac{\log n}{r^{2d}}\right)$ (such $z$ must exist, as there is a logarithmic amount of crossings of dense cubes in each strip), we apply again the triangle inequality and we are done.
To generalize to different slices, we proceed as in the last step of the proof of Theorem 5.1, were we saw that we may go from one slice to another in a (fixed) maximum of steps. Subsequent applications of the triangle inequality are enough to obtain the desired result.
We thus finished proving that a.a.s. the effective resistance between any two vertices that belong to dense cubes that belong to the giant component of dense cubes is $O\left(\frac{\log n}{r^{2d}}\right)$. ∎
Lemma 6.2.
Let $r_{0}$ be a large enough constant, and let $r_{0}\leq r\leq(1-\varepsilon)r_{c}$. Then
$\tau_{cov}(L_{1}(G))=O(n\log^{2}n)$ (6.4)
Proof.
It is a standard topological fact that every (maximal) connected component of cubes, which might contain dense and non-dense cubes and that has empty intersection with the giant component of dense cubes, has a diagonally connected outer boundary of non-dense cubes. As there are at most $(3^{d}-1)^{L}$ diagonally connected paths of length $L$ when the starting vertex is fixed, and since the probability that a cube is non-dense is $e^{-\kappa r^{d}}$ for some $\kappa=\kappa(\beta)>0$, the probability that there exists a diagonally connected path of non-dense cubes of length greater or equal than $L$ is, by a union bound over all $(1+o(1))n$ starting vertices, at most
$(1+o(1))ne^{-L\kappa r^{d}}(3^{d}-1)^{L},$
which tends to 0 for $L:=c\frac{\log n}{r^{d}}$ with $c>0$ large enough. Therefore, since for any $M>0$, every connected component (in the renormalized lattice of cubes $\mathrm{Q}$) of volume $M^{d}$, by a standard isoperimetric inequality has to have an outer boundary of length at least changed this, it was only for 2 dimension $CM^{d-1}$ for some $C>0$, a.a.s. every component different from the giant of the associated Bernoulli site percolation has volume at most $O(L^{d/(d-1)})$. In other words, for every site, at distance $L$ one finds a site belonging to the giant component (of the renormalized lattice of dense cubes), and for any vertex $x$ of the underlying random geometric graph, at Euclidean distance at most $O(Lr/\mu)=O(\log n/r^{d-1})$ we find a vertex $y$ belonging to a cube of the giant component of the Bernoulli site percolation.
Then, by Corollary 2.12, a.a.s. we have $d_{G}(x,y)=O\left(\log n/r^{d}\right)$, and this implies that $\mathcal{R}(x\longleftrightarrow y)=O(\log n/r^{d})$. If $z\in G$ is another vertex that belongs to some dense cube in the backbone $\mathcal{C}^{site}_{\mathrm{Q}}(n)$ then by  Theorem 6.1 we have $\mathcal{R}(y\longleftrightarrow z)=O(\log n/r^{2d})$, and hence by the triangle inequality, also
$\mathcal{R}(x\longleftrightarrow z)=O(\log n/r^{d})+O(\log n/r^{2d})=O(\log n/r^{d}).$
Since by Lemma 2.10, the number of edges of the giant component satisfies $|E|=\Theta(nr^{d})$ a.a.s., by Theorem 2.6, we have that
$\tau_{cov}(L_{1}(G))=O(n\log^{2}n),$
and the lemma follows. ∎
Proof of Theorem 1.1. By combining Theorem 4.2 and Theorem 5.1 with Lemma 6.2, Theorem 1.1 follows immediately. ∎
7. Upper bound right above the threshold for connectivity
Perhaps surprisingly, for $\mathcal{G}(n,r,2)$, no upper bound of $O(n\log n)$ on the cover time of the simple random walk had been known right above the connectivity threshold. For $d\geq 3$, in the already mentioned paper of Cooper and Frieze [12] the result was obtained, but their result does not apply for $d=2$. As mentioned in the introduction, for $d=2$, in [5] the authors found the asymptotic order of the cover time only for a radius that was by a constant factor larger than the connectivity threshold. In this section, we close the gap left by the papers for $d=2$. We note that it is enough to prove the upper bound, as the lower bound is immediate by the general lower bound on cover times for connected graphs by Feige [19]. Our main result will therefore be the following theorem (since our result is only new for $2$ dimensions, we state it only for 2 dimensions, although it holds more generally):
Theorem 7.1.
Let $\varepsilon>0$ arbitrarily small and consider $G\in\mathcal{G}(n,r,2)$. Suppose that $r^{2}=(1+\varepsilon)\frac{\log n}{\pi}$. Then, a.a.s, for any two vertices $x,y\in G\in\mathcal{G}(n,r,2)$,
$\mathcal{R}(x\longleftrightarrow y)=O\left(\frac{1}{\log n}\right).$ (7.1)
Proof of Theorem 1.2 assuming Theorem 7.1: By Lemma 2.10, a.a.s., $|E|=\Theta(nr^{d})=\Theta(n\log n)$. Theorem 1.2 then follows immediately by Theorem 2.6. ∎
It thus remains to prove Theorem 7.1. The approach of the proof is similar to the one in Section 6: first we prove that the upper bound on the effective resistance is true for any pair of vertices inside the backbone of dense cubes, and in a second step we then prove that the resistance between any two vertices is of the desired order. We now proceed to the details:
Proof of Theorem 7.1: By Theorem 6.1, the result holds for any $x,y$ with $x\in\mathrm{Q_{i}},y\in\mathrm{Q_{j}}$, and $\mathrm{Q_{i}},\mathrm{Q_{j}}\in\mathcal{C}^{site}_{\mathrm{Q}}(n)$. It thus remains to show that the effective resistance between any vertex $x\in G$ that is in a cube, being dense or non-dense, that is not in the giant component of the auxiliary Bernoulli site process, and some vertex $y\in G$, belonging to a dense cube in the giant component of the auxiliary Bernoulli site process is at most $O\left(\frac{1}{\log n}\right)$: by the triangle inequality (Lemma 2.4) together with Theorem 6.1, this will then show that for any pair of vertices $x,y\in G$, we have $\mathcal{R}(x\longleftrightarrow y)=O\left(\frac{1}{\log n}\right)$.
In order to show this bound on the effective resistance, it suffices to show that for every vertex $x$ we find $y$ belonging to the giant of the auxiliary Bernoulli site process together with a set $\mathcal{P}_{x,y}$ of $\Omega(\log n)$ internal vertex-disjoint paths between $x$ and $y$ of constant length. Indeed, in this case one can split evenly the unit flow emanating from $x$ between all these paths, such that the flow on each such edge of each path is $O(1/\log n)$, and the resistance, by Definition 2.1, thus satisfies
$\mathcal{R}(x\longleftrightarrow y)=\sum_{e\in P\in\mathcal{P}_{x,y}}O(1/\log^{2}n)=O\left(\frac{1}{\log n}\right).$
We will follow this plan in detail now. We first assume that $x$ is at distance at least $3C\sqrt{\log n}$ from all boundaries of $\Lambda_{n}$, where $C$ is a sufficiently large constant.
Recall that by the proof of Theorem 6.1, for every $x\in G$ there exists $y\in G$ at distance $rL/\sqrt{5}=O(\log n)$, with $y$ belonging to a dense cube of the giant component of the auxiliary Bernoulli process (and therefore also $y\in L_{1}(G)$). We thus may find, an axis-parallel square $S$ of side length $3C\sqrt{\log n}$ containing both $x$ and $y$, such that both $x$ and $y$ are at distance at least say $C\sqrt{\log n}$ from the boundary of $S$, and such that the left bottom corner of $S$ is an integer coordinate. We now show that we cannot separate $x$ from $y$ by removing less than $\varepsilon^{\prime}\log n$ many vertices inside $S$ (and therefore we cannot separate $x$ and $y$ in $G$ either), for some $0<\varepsilon^{\prime}<\varepsilon$. Since this will hold a.a.s. for all $x,y$ simultaneously (with $S$ changing according to positions of $x,y$), we conclude that the graph is $\varepsilon^{\prime}\log n$-connected. By Menger’s theorem, for every pair of vertices there exist $\varepsilon^{\prime}\log n$ internally vertex-disjoint paths inside $S$ between any pair of vertices, and we will show then that they can be chosen of constant length.
To do so, we adapt ideas from a classic $k$-connectivity argument on random geometric graphs (it is less precise than the proof of Theorem 1.1, in particular Proposition 5.2 in Penrose’s paper [34], but some ideas are adapted), applied with $k=\varepsilon^{\prime}\log n$. First, we show that after removing less than $\varepsilon^{\prime}\log n$ many vertices, $G$ a.a.s. does not have two components $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ both of Euclidean diameter larger than $C\sqrt{\log n}/5$ ($C$ the same as before). In fact, we show that this cannot happen inside $S$. Indeed, if this were the case, then consider the cells of the tessellation with cells of sidelength $r/\sqrt{20}$ forming the exterior boundary of the smaller of the two components, say $\mathcal{C}_{1}$ (formally, this exterior boundary is defined as the exterior boundary of the component which contains cells that either contain a vertex of $\mathcal{C}_{1}$ or contain an edge of $\mathcal{C}_{1}$, where an edge is drawn is a straight line). It is a standard topological fact that in $\mathbb{R}^{2}$ the outer boundary is diagonally connected, and its restriction to $\Lambda_{n}$ contains at least a quarter of this boundary. The number of cells of this outer boundary inside $\Lambda_{n}$ is least $(C\sqrt{\log n}/20)/(r/\sqrt{20})\geq C/4$. In fact, by choice of $S$, this lower bound holds also for the longest connected part of the exterior boundary inside $S$. Observe that the maximal Euclidean distance between two vertices in any two diagonally adjacent cells is at most $\sqrt{3r^{2}/20}\leq r/2$, and hence, if there were any vertex in a cell of the exterior boundary, such a vertex would be connected by an edge to a vertex in $\mathcal{C}_{1}$ (indeed, if the neighboring square was hitting only an edge of $\mathcal{C}_{1}$, its distance to the edge were at most $r/2$, and from there to the closer endpoint of the edge at most $r/2$, so in total at most $r$). Now, denote by $X$ the number of vertices in a fixed set of $C/4$ cells of radius $r/\sqrt{20}$. We have by standard Chernoff bounds for Poisson random variables
$\Pr(X\leq\varepsilon^{\prime}\log n)=\Pr(\mathrm{Poi}(Cr^{2}/80)\leq\varepsilon^{\prime}\log n)\leq(eC/(\pi\varepsilon^{\prime}80))^{\varepsilon^{\prime}\log n}e^{-Cr^{2}/80}\leq e^{-Cr^{2}/200},$
assuming $C$ large enough and $\varepsilon^{\prime}<\varepsilon$. The number of such lattice animals is at most $O(n8^{C/4})$, and thus, by Chernoff bounds together with a union bound over all such animals, the probability of having in one of them less than $\varepsilon^{\prime}\log n$ many vertices is at most $O(n8^{C/4}e^{-r^{2}C/200})=O(n8^{C/4}e^{-(1+\varepsilon)C\log n/(200\pi)})=o(n^{-2})$, where the equality follows by choosing $C$ large enough. Hence, for $C$ large, by taking a union bound over all pairs of vertices, a.a.s. such two components cannot exist.
Next, if after removing less than $\varepsilon^{\prime}\log n$ vertices there would exist a connected component $\mathcal{C}_{1}$ of Euclidean diameter at most $C\sqrt{\log n}/5$ inside $S$, this component a.a.s. has $O(\log n)$ vertices. If the Euclidean diameter of such a component $\mathcal{C}_{1}$ were less than $\varepsilon^{\prime\prime}\sqrt{\log n}$ (for $\varepsilon^{\prime\prime}<\varepsilon^{\prime}<\varepsilon$ small enough; we define the Euclidean diameter of an isolated vertex to be zero and include this in this case), then note that the probability to have at most $\varepsilon^{\prime}\log n$ points in the annulus obtained by taking the difference of the ball of outer radius $r$ and inner radius $\varepsilon^{\prime\prime}\sqrt{\log n/\pi}$ centered around any of the vertices in $\mathcal{C}_{1}$, is, once again by Chernoff bounds for Poisson variables, at most
$O(\log n)\Pr(\mathrm{Poi}((1+\varepsilon-\varepsilon^{\prime\prime})\log n)\leq\varepsilon^{\prime}\log n)=(e(1+\varepsilon-\varepsilon^{\prime\prime})/\varepsilon^{\prime})^{\varepsilon^{\prime}\log n}e^{-(1+\varepsilon-\varepsilon^{\prime\prime}+o(1))\log n}\leq n^{-1-\varepsilon^{\prime\prime\prime}},$
for some $\varepsilon^{\prime\prime\prime}>0$, where we assumed $\varepsilon^{\prime\prime}<\varepsilon^{\prime}$. By a union bound over all at most $n(1+o(1))$ components such a component a.a.s. does not exist.
Otherwise, if the Euclidean diameter were at least $\varepsilon^{\prime\prime}\sqrt{\log n/\pi}$ (but still at most $C\sqrt{\log n}/5$, and therefore still completely inside $S$), then observe that the smallest axis-parallel rectangle $R$ containing all vertices of such a component has to have without loss of generality horizontal length at least $(\varepsilon^{\prime\prime}\sqrt{\log n/\pi})/2$. A.s., such rectangle (of area $O(\log n)$) is defined by at most $4$ points, and hence, for every fixed $S$, there are at most $O(\log^{4}n)$ choices for the rectangle. Therefore, the halfcircle of radius $r$ to the left of the leftmost point of $R$ as well as the halfcircle to the right of the rightmost point of $R$ have to contain at most $\varepsilon^{\prime}\log n$ points (otherwise by removing less than $\varepsilon^{\prime}\log n$ vertices $\mathcal{C}_{1}$ is not disconnected). The probability for this to happen is, by a union bound over all $O(n)$ choices for $S$ and thus in total over all at most a.a.s. $O(n\log^{4}n)$ choices of rectangles, by another Chernoff bound for Poisson random variables, at most
$O(n\log^{4}n)\Pr(\mathrm{Poi}((1+\varepsilon)\log n)\leq\varepsilon^{\prime}\log n)\leq(e(1+\varepsilon)/\varepsilon^{\prime})^{\varepsilon^{\prime}\log n}n^{-\varepsilon+o(1)}=o(1),$
where we again assumed $\varepsilon^{\prime}<\varepsilon$. Hence, a.a.s. such component also cannot exist, and for any $x\in G$ we find inside $S$ the desired vertex $y$.
We now sketch the modifications when $x$ is close to the boundary: suppose first that $x$ is at distance at most $3C\sqrt{\log n}$ from exactly one boundary of $\Lambda_{n}$. In case of having two components $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ of diameter larger than $C\sqrt{\log n}/5$ after removing less than $\varepsilon^{\prime}\log n$ vertices (one of them containing $x$), replacing $S$ from before by $S\cap\Lambda_{n}$, the same argument as before shows that this cannot happen a.a.s. Similarly, a component $\mathcal{C}_{1}$ containing $x$ of diameter less than $\varepsilon^{\prime\prime}\sqrt{\log n}$ cannot happen, as the semiannulus around $x$ obtained by the difference of the ball of radius $r$ around $x$ and the one of radius $\varepsilon^{\prime\prime}\sqrt{\log n}$ cannot be empty a.a.s. (the union bound again is only over all $n^{1/2+o(1)}$ many vertices). Also, if the Euclidean diameter is of size at most $C\sqrt{\log n}/5$, but larger than $\varepsilon^{\prime\prime}\sqrt{\log n}$, a semicircle of radius $r$ plus an extra area of size $\Theta(r)$ would have to be empty, and a.a.s. by a union bound over all $n^{1/2+o(1)}$ vertices this cannot happen.
In case $x$ is at distance at most $3C\sqrt{\log n}$ from two boundaries of $\Lambda_{n}$, note that there are only $n^{o(1)}$ such vertices a.a.s., and the adaptations of the other arguments are also analogous.
Hence, as a conclusion, we see that $G$ restricted to $S\cap\Lambda_{n}$ is a.a.s. $\varepsilon^{\prime}\log n$-connected. By Menger’s theorem, for any $x,y$ there exist thus $\varepsilon^{\prime}$ internally vertex-disjoint paths between $x$ and $y$ inside $S$. Moreover, since these paths are inside $S$, one may shorten these to make them paths of constant length by making it an induced path: for each path from $x$ to $y$, we shorten it by maintaining only the first and last vertex of the path passing through any cube $\mathrm{Q_{i}}$ inside $S$ (the path in between is replaced by a direct edge between these vertices), and hence its length is at most constant. The theorem follows.
8. Concluding remarks and future work
We showed that the cover time undergoes a jump at the threshold of connectivity, decreasing from $\Theta(n\log^{2}n)$ to $\Theta(n\log n)$ a.a.s. It is therefore natural to ask what happens right at the threshold of connectivity: assume then that $r^{d}=(\log n+f(n))/V_{d}$, where $f(n)$ may be positive or negative, and $|f(n)|=o(\log n)$. If $f(n)$ is negative and $|f(n)|=\omega(1)$, the same argument as in Case 2 of Section 4, can be used to show that the number of vertices of degree $1$ inside the giant component is a.a.s. $\Omega(nr^{d}V_{d}e^{-r^{d}V_{d}})=\Omega(e^{-f(n)})$ (which is positive since $f(n)$ is negative). Hence, we obtain a set $V^{\prime}$ of those vertices of degree 1 vertices $x,y$ for which $\log(|V^{\prime}|)=-f(n)$, and any pair of vertices of $x,y\in V^{\prime}$ satisfies $\mathcal{R}(x\longleftrightarrow y)\geq 1$. Thus, by Theorem 2.7, a.a.s. $\tau_{{\rm cov}}\geq\gamma n\log n|f(n)|$. We do not know for which values of $f(n)$ this lower bound is tight. On the other hand, the upper bound of Theorem 6.1 shows that between any two vertices $x,y$ belonging to a dense cube of the auxiliary giant component (as defined there), $\mathcal{R}(x\longleftrightarrow y)=O(1/\log n)$. The value of the cover time is thus determined by the maximal resistance between pairs of vertices, one of which being outside the giant. It is subject of further work to obtain results on these pairs.
More generally, in the same spirit as the very precise works on the cover time for the random walk on $\mathbb{Z}^{d}$, one could ask for more precise estimates on the cover time for any $r\leq r_{c}$. Since we do not have very precise estimates on the Green’s function (this being essential for the calculations there) and we neither know of an isomorphism translating $\mathcal{G}(n,r,2)$ into a problem on Gaussian free fields, this seems to be out of reach with the current techniques.